### 28 Dec 2023 v1

# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

Abstract
Emergence and causality are two fundamental concepts for understanding complex systems. They are interconnected. On one hand, emergence refers to the phenomenon where macroscopic properties cannot be solely attributed to the cause of individual properties. On the other hand, causality can exhibit emergence, meaning that new causal laws may arise as we increase the level of abstraction. Causal emergence theory aims to bridge these two concepts and even employs measures of causality to quantify emergence. This paper provides a comprehensive review of recent advancements in quantitative theories and applications of causal emergence. Two key problems are addressed: quantifying causal emergence and identifying it in data. Addressing the latter requires the use of machine learning techniques, thus establishing a connection between causal emergence and artificial intelligence. We highlighted that the architectures used for identifying causal emergence are shared by causal representation learning, causal model abstraction, and world model-based reinforcement learning. Consequently, progress in any of these areas can benefit the others. Potential applications and future perspectives are also discussed in the final section of the review.
keywords:
Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model
[…]
1 Introduction
Economic growth, environmental protection, sustainable development, the global climate crisis, social inequality, and many other issues are all intertwined with complex systems Ledford (2015); Mensah et al. (2018). Therefore, gaining a deep understanding of how complex systems operate, evolve, grow, stabilize, and collapse is of paramount importance. However, this task is exceptionally challenging due to the fact that complex systems consist of diverse and heterogeneous agents that interact through complex nonlinear relationships Bar-Yam (2002). Moreover, they all exhibit emergent phenomena, which are highly common in complex systems but carry a sense of mystery Holland (2000).
How did the first living cell emerge from the collisions between various molecules in the early Earth’s environment Artime and De Domenico (2022)? How does the cognitive concept of "I" emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)? How do large neural language models suddenly exhibit emergent abilities Wei et al. (2022)? These fundamental questions revolve around the concept of emergence in complex living, cognitive, and artificial systems. Emergence refers to the phenomenon where macroscopic properties and phenomena that cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000). This presents a formidable challenge to the traditional reductionist perspective, while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
However, as elucidated by Bedau’s theory of weak emergence Bedau (1997), many emergent phenomena can be comprehended through the interactions among the individuals within the system Holland (2000). Complex systems, in fact, consist of extensive networks of interacting components Albert and Barabási (2002). Within these networks, even a minor cause, such as a perturbation of a single unit, can propagate through the interconnected network, resulting in a collective effect. The phenomenon described is commonly known as the butterfly effect Wikipedia (2023a), which provides an explanation for the occurrence of emergence. On the other hand, emergent properties, such as homeostasis Wikipedia (2023b), can stabilize the system itself, thereby preserving the locality of causal effects and preventing the observation of macro-level effects. These phenomena demonstrate that complex systems achieve interactions through causal laws, where numerous local causal laws form interconnected causal networks as a whole. And this whole possesses its unique causal characteristics.
Causality, or causation, refers to the connection between a cause and its resulting effect Granger (1969); Pearl et al. (2000); Pearl (2009). It describes the phenomenon in which an event, known as the cause, leads to another event, known as the effect. Traditional studies of causality have typically focused on the causal relationship between two or a few variables. However, the unique characteristics of causality in complex systems present new challenges to classical causal science due to the vast number of variables involved and the presence of emergent phenomena. In complex systems, it is possible for one cause to have multiple effects, and conversely, one effect may be influenced by a multitude of causes. Furthermore, in complex systems, causality often exhibits cross-level properties, which are closely associated with emergence.
Emergence and causation interconnect with each other. On the one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956). On the other hand, emergent properties may have causal effects on individuals in complex systems Kim (1992); Holland (2000). For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market, while the price may also give feedback to the market: it may affect the decision making of each individual.
Furthermore, we can understand emergence through the perspective of causation. What does emergence mean is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005). Thus, emergent properties or phenomena lose their direct explanations as usual, but may be attributed to the causes on the macro-level as pointed out by Hoel et al. (2013). Therefore, new causalities can be observed on larger scales.
In conclusion, gaining a deep understanding of emergence is crucial in the field of complex system studies. Specifically, the development of a quantitative theory of emergence is on the verge of emerging. Such a theory holds the potential to address significant challenges, including the origins of life, the emergence of novel capabilities in large neural network models, and the potential for intelligence, consciousness, and free will to arise in artificial systems. Causality not only exhibits a profound connection with emergence but is also considered by many researchers as one of the most crucial perspectives for quantitatively comprehending emergence Hoel et al. (2013); Crutchfield and Young (1989); Fromm (2005); Seth (2008).
Two primary challenges take precedence in understanding emergence from a causal perspective. The first is establishing a quantitative definition of emergence, while the second involves identifying emergent behaviors or phenomena through data analysis.
To address the first challenge, two prominent quantitative theories of emergence have emerged in the past decade. The first is Erik Hoel’s theory of causal emergence Hoel et al. (2013), while the second is Rosas’ theory of emergence based on partial information decomposition Rosas et al. (2020).
Hoel’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains. It employs the concept of effective information (EI) to quantify the extent of causal influence within Markov chains and enables comparisons of EI values across different scales Hoel et al. (2013); Hoel (2017). Causal emergence is defined by the difference of EI values between macro-level and micro-level. Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than the micro-level. Hoel and other researchers further extended measures of effective information and causal emergence for dynamical systems with continuous variables in Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020). Other measures are also possible to quantify causal effects and thereafter causal emergence. In Comolatti and Hoel (2022), Hoel systematically compared a bunch of causal effect measures and concluded that causal emergence is independent on the selection of the measure.
However, it is essential to establish a coarse-graining strategy beforehand in Hoel’s theory of causal emergence. Alternatively, the strategy can be derived by maximizing the effective information (EI) Hoel et al. (2013). However, this task becomes challenging for large-scale systems due to the computational complexity involved. To address these problems, Rosas introduced a new quantitative definition of causal emergence Rosas et al. (2020) that does not depend on coarse-graining methods, drawing from Partial Information Decomposition (PID) related theory. PID is an approach developed by Williams et al., which seeks to decompose the mutual information between a target and source variables into non-overlapping information atoms: unique, redundant, and synergistic information Williams and Beer (2010). Based on this groundwork, Rosas further developed the concept and introduced a theory called $\phi$ID to decompose the mutual information between multiple target and source variables Mediano et al. (2019). This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between source and target variables based on the inherent characteristics of the system.
The second challenge pertains to the identification of emergence from data. In an effort to tackle this issue, Rosas et al. have derived a numerical method Rosas et al. (2020). However, it is important to acknowledge that this method offers only a sufficient condition for emergence and is an approximate approach. Another limitation is a coarse-grained macro-state variable should be given beforehand to apply this method. Hence, there is a need for the development of new methods. In a recent study by Zhang et al. Zhang and Liu (2022), a machine learning framework named Neural Information Squeezer (NIS) was introduced to tackle the challenge of identifying causal emergence using Hoel’s definition. Remarkably, the NIS neural network, functioning as a "machine observer" equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data. This aligns with the viewpoint suggested by Eric Bonabeau and Jean-Louis Dessalles Bonabeau and Dessalles (2011) that all emergent properties and phenomena are observer-dependent, and the epistemological nature of emergent causal powers, including downward causation Kim (1992); Crutchfield and Young (1989). However, employing a machine observer offers additional advantages in identifying emergence from data, as it helps avoid subjective judgments.
Furthermore, we find that causal emergence identification and the recent development of artificial intelligence can have deeper connections through the bridge of NIS and its extended frameworks Yang et al. (2023). First, the architecture of NIS, that is, the “encoder - dynamics learner - decoder” framework, is similar to causal representation learning Schölkopf et al. (2021) in artificial intelligence where dynamics learner is replaced as causal mechanisms. Therefore, there is a deep connection between the two emerging fields: causal emergence and causal representation learning. The former tries to characterise and understand the emergent causality in a complex system and the latter tries to learn the latent causal structures and mechanisms from observational data. They are connected by the problem of causal emergence identification. That is to identify causal emergence from data is essentially equivalent to learning the latent causality inherent in the data.
Second, there are more connections between causal emergence and other areas in artificial intelligence. For example, causal model abstraction also discussed the interplay between causality and the abstraction (coarse-graining) of causal models Beckers et al. (2020). In reinforcement learning based on world models Ha and Schmidhuber (2018), the similar framework, “encoder - dynamics learner - decoder”, is also used. Thus, the similarity and the shared features between the two fields can help us to lend the ideas and techniques in one field to the other field. For instance, EI may be used to measure the causal effect of neural networks Marrow et al. (2020), and the techniques in causal representation learning can be employed to detect and identify causal emergence. Zhang and Liu (2022).
This paper tries to review the latest works on the quantitative theory and applications of causal emergence, and its relative works. We also discuss its connections with artificial intelligence and complex systems. In the second part, we introduce the background of causal emergence, particularly on the connections between causation and emergence in complex systems. The third part introduces several quantitative theoretic frameworks including: Crutchfield et al’s computational mechanics Crutchfield and Young (1989), Seth et al’s granger causal emergence Seth (2008), Hoel et al’s causal emergence theory and Rosas et al’ theory of emergence based on information decomposition. Several related conceptions are discussed, such as coarse-graining strategy, measure of effective information, and partial information decomposition. We also compare these two theories from several aspects. The forth part discusses the problem of causal emergence identification and then introduces the connection with artificial intelligence. Causal representation learning Schölkopf et al. (2021), abstraction of causal models Beckers et al. (2020), and world models in reinforcement learning Ha and Schmidhuber (2018) are introduced and the relations with causal emergence identification are discussed. Finally, we discuss the connections between causal emergence and other important topics in complex systems such as the origins of life, consciousness, etc.
2 Background: from Causality to Emergence
2.1 Causality
Causality, which is a fundamental concept in many fields, including philosophy, natural science, and social science, refers to the relationship between a pair of events, where the first event (cause) can influence the second one (effect) Granger (1969); Pearl et al. (2000); Pearl (2009). The relationship between cause and effect can exhibit either deterministic or probabilistic characteristics. In deterministic causality, the cause will always produce the same effect, while in probabilistic causality, the cause will only produce the effect with a certain probability.
In most physical systems governed by differential equations or Markovian dynamics, whether deterministic or probabilistic, causality is inherent. This arises from the fact that manipulating a variable within the system can result in observable changes in other variables, albeit with probabilistic outcomes. While, one physical mechanism may correspond several possible causal structures as pointed by Y. Iwasaki and H. A. Simon Iwasaki and Simon (1994) because the causal graphs describe local causal interactions while mechanisms are the global constraints for all variables.
Figure 1: The causal models (left column) and the corresponding causal hierarchy (right column) and .
However, if we do not know the physical mechanism of a system, how can we model the causality behind the system? There are several parallel theoretical frameworks to do this. For example, Judea Pearl tries probabilistic graphic models (e.g. Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al. (2000); Pearl (2009). In which nodes represent random variables, acyclic links represent causal interactions. Three different causal hierarchies are distinguished and quantified in different models. The first level of causality is association, i.e., the correlation between variables, which can be modelled by Bayesian networks Sucar (2015). The second level is intervention, which can be characterised by causal graph models Pearl (2009). The difference between a causal graph and a Bayesian network is the former allows defining a special operator which is called "do" operator to simulate the interventions done by the experimenter. The last level is courterfactual, which can be quantified by structural causal models Pearl et al. (2000). The distinguishing characteristic of a structural causal model is that all interactions within a causal graph can be described by a set of deterministic functions that involve the input variables and unknown components, thereby introducing uncertainty. Counterfactual causal inferences, that is the inferences under an imaginary scenario, can be conducted only on structural causal models. The different types of causal models and the corresponding causal hierarchy in Pearl’s theory are shown in Figure 1.
Another quantitative framework for causal inference is Donald B Rubin’s "potential outcome" theory which is directly defined on the level of counterfactual Rubin (2005). According to this theory, each individual has two potential outcomes: the outcome that would be observed when individuals receive the treatment and the outcome that would be observed when they do not receive the treatment. Rubin’s theory suggests that the causal effect of a treatment can be estimated by comparing the difference between the observed outcome and the unobserved potential outcome for each individual.
Researchers also developed a bundle of methods, which are named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data. For example, constraint-based algorithms Spirtes et al. (2000) and score-based algorithms Chickering (2002) for causal discovery are two typical representations . Causal discovery has become a hot interdisciplinary area in the middle of machine learning, causal inference, and statistics.
Due to the uncertainty and ambiguity behind the discovered causal relations, measuring the degree of causal effect between two variables is another important problem. Numerous independent historical studies have tackled the issue of causation measurement. These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009), etc. The effective information (EI) in the framework of causal emergence is also an indicator to measure causal effect. All these measures can almost be decomposed into two terms which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
In summary, causality is a fascinating inter-disciplinary topic across for several subjects. Quantitative framework and measures have been developed in last decades.
2.2 Emergence
Emergence, known as "the whole is greater than the sum of parts" Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro and macro levels of organization. Therefore, the scale of a system (micro or macro) and the cross-level interactions must be considered when we discuss emergence. This may increase the complexity of study for emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al. (2019); Ross et al. (1908); Holland (1996)
Emergence has deep connections with causation Fromm (2005); Bedau (1997); Bedau and Humphreys (2008); Bedau (2002). First, emergence emphasizes the new properties arising from the interactions between the components in the system, rather than the components themselves. Consequently, all emergent properties and processes at the macro-level can, in principle, be comprehended as the causal effects resulting from interactions between individuals at the micro-level. This is what the philosophical notion “supervenience” tries to describe.
However, it does not mean that all the emergent phenomena on macro-level can be easily attributed to micro-level individuals and their interactions. Two major reasons can account for this. The first reason is due to the principle of computational irreducibility proposed by Wolfram Wolfram et al. (2002) which claims that although we know the emergent properties are out of a set of simple rules, we can not find a short-cut way in principle to predict the results except to implement the rules. For example, complex behaviors (eg. herding and aligning) of the Boid model are the results of simple rules, but these behaviors can not be reduced to individual behaviors in a simple way except to simulate it Reynolds (1987).
The second reason is that emergent behaviors may be determined by other emergent properties on macro-level Hoel et al. (2013); Rosas et al. (2020). For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because the severe lack of fossil oil may cause inflation. Therefore, the cause of an emergent property (rice price) is another emergent macro-level coarse-grained variable and it means new causation on macro-level coarse-grained variables can be observed as pointed by the theory of causal emergence. And, interestingly, the causation in different levels must be consistent to explain the micro-level event of the price of rice. Otherwise, double causation Merton (1973) fallacy may appear. This consistency between different scale causality is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
In order to better understand emergence, Bedau et al Bedau (1997, 2002) classify emergence into three categories according to the causal interactions between micro and macro, which are nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996) and strong emergence O’Connor (1994); Kim (1992).
In all these categories, the notion of nominal emergence is least controversial, which can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994). For example, the patterns which are composed of many pixels on a screen are nominal emergence. We can say nominal emergent patterns as “supervenient” because all the macro-level properties are from individuals.
While weak emergence refers to macro-level properties or processes that can be derived through interactions between individual components but in a complicated way such that they are hardly reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996). For the weak emergent patterns or processes, the causes may come from both micro-level and macro-level, therefore, emergent causation may coexist with micro-level causation.
There are many debates on strong emergence, which refers to properties or processes on macro-level that can not be reduced to micro-level properties in principle, including the interactions of individuals O’Connor (1994); Kim (1992). Thus, the collective behaviors of boids are not strong emergent because they are the results of interactions between Boid model. This notion is controversial because it rejects any mechanistic explanations for strong emergent properties, which can not be explained by micro-level variables. Second it raises some long standing philosophical debates about “causal fundamentalism”  Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
Jochen Fromm further explains strong emergence as the causal effect of downward causation Fromm (2005). Consider a system contains three-different scales: micro, meso, and macro. Downward causation refers to the causal power from the macro-level to the messo-level or from the messo-level to the micro-level. Thereafter, although the strong emergent properties or processes on the meso-level are supervenient to micro-level properties and interactions, their causes are from the macro-level and therefore a second order patterns or processes. However, there are many debates on the notion of downward causation itself, e.g.  Yurchenko (2023); Bedau and Humphreys (2008).
Figure 2: Various types of causation that exist within multi-scale complex systems. In this figure, solid arrows represent common causation, which is generally accepted without much controversy. Dotted arrowed lines indicate a form of “causation” driven by supervenience, while dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
In summary, there are three types of causation based on their cross-level properties, which are associated with emergence, as depicted in Figure 2:
1. Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
2. Intra-level causation: This refers to causal effects occurring within the same scale or level.
3. Downward causation: Here, macro-level properties have influences on micro-level properties.
In Figure 2, the black solid lines represent causation without controversy. However, according to a recent paper by Yurchenko  Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality, but rather reasons or explanations. Yurchenko introduces the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers. Instead, they serve as reasons or explanations from the perspective of an observer. While, Yurchenko accepts the intra-level causation in the meso- or macro-level as “real” as the ones in micro-level. He even coined the term “causal equivalent principle” to represent the idea that all the intra-level causation except cross-leveled ones should be accepted and are equivalent each other.
However, other scholars think that the emergent causation, the intra-level causation in the meso- or macro- level, depends on observers Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022). There are much controversy on emergent cross-leveled causation.
3 Quantifying Emergence by Causality
Since causality and emergence have a strong connection, and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence. In this section, we will review several frameworks to quantify multi-scale causality and emergence.
3.1 Early Related Works
Before the theory of causal emergence was proposed by Hoel et al, there were some related works that have introduced very similar ideas as causal emergence theory. For example, Crutchfield et al’s computational mechanics theory considered causal states which are the partitions of the state space and may have good predictions. While, Seth et al proposed G-causality to quantify emergence by using Granger’s causality. We will introduce them in detail.
3.2 Computational Mechanics
The theory of computational mechanics proposed by Crutchfield, Shalizi, and Feldman et al. tries to formulate this kind of emergent causation in a quantative framework Crutchfield and Young (1989). In some sense, computational mechanics can be understood as the inverse of statistical mechanics because what statistical mechanics do is to derive macro-level consequences from micro-dynamics, and the inverse process done by computational mechanics is to construct a minimal causal model from the observations of a stochastic process that can generate the observed time series.
Let’s assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$. We can divide it into two segments beginning from time step $t$: the history before $t$ which is denoted as $\overleftarrow{s_{t}}$ and the future after $t$ which is denoted as $\overrightarrow{s_{t}}$. If the process is stationary, we can drop the denotation of $t$. Thus, all the possible histories $\overleftarrow{s}$ form a set which is denoted as $\overleftarrow{S}$ and all the futures $\overrightarrow{s}$ form a set $\overrightarrow{S}$.
We aim to establish a model that can reconstruct and predict observed random sequences, with a higher accuracy being desirable. However, the randomness of the sequences prevents us from achieving a perfect reconstruction unless we record every randomly occurring character. This would make the model excessively long. To preserve useful information as concisely as possible, we need a coarse-grained mapping that captures the ordered structure in the random sequences, known as patterns Crutchfield (1994). We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets which form a set $\mathcal{R}$. Any subset in $R\in\mathcal{R}$ is a state. We define a function from histories to states, $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$. Thus $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
For a set of states $\mathcal{R}$, we can measure its simplicity using a complexity metric. Intuitively, the larger the cardinality of $\mathcal{R}$, the more complex it is. Additionally, we need to consider its distribution. For example, if one state occurs frequently while others occur rarely, it is less complex compared to a situation with a uniform distribution. Therefore, we can define the statistical complexity $C_{\mu}$ of a set of states using Shannon entropy Shalizi (2001):
$C_{\mu}(\mathcal{R})\triangleq-\sum_{\rho\in\mathcal{R}}P(\mathcal{R}=\rho)\log_{2}P(\mathcal{R}=\rho).$ (1)
When constructing predictive models using a set of states, the statistical complexity refers to the size of the model.
What kind of state set can achieve the best balance between predictiveness and parsimony? We can introduce an equivalent relationship which is called causal equivalence Crutchfield (1994). Concretely, we say $\overleftarrow{s}$ and $\overleftarrow{s}^{\prime}$ are causally equivalent if and only if:
$P(\overrightarrow{s}|\overleftarrow{s})=P(\overrightarrow{s}|\overleftarrow{s}^{\prime})$
This equivalent relationship can partition all the histories into equivalent classes, and they are defined as causal states. We denote all the causal states of a history $\overleftarrow{s}$ as: $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}}$ which is a subset of the histories.
For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that make the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
$T^{(s)}_{ij}\equiv\text{P}(\mathcal{S}^{\prime}=\mathcal{S}_{j},\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s|\mathcal{S}=\mathcal{S}_{i}),$
where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and the notation $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$ in which the first symbol emitted is $s$111we use slightly different symbols compared with reference  Shalizi (2001). Hence, by combining: $\mathcal{S^{\prime}}=\mathcal{S}_{j}$ and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$, we ensure that all subsequent causal states will have the identical initial emitted symbol $s$. We denote the set $\{T^{(s)}_{ij}:s\in\mathcal{A}\}$ by $T$. And the definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$) and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
Up to now we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories. The remaining task is to show that it is in some sense optimal. It has the important properties of being minimally predictive, maximally statistically complex, and minimally stochastic Shalizi (2001).
We shall compare the causal state $\mathcal{S}$ with any state $R\in 2^{\overleftarrow{\mathcal{S}}}$, to show that for the purpose of predicting the future, the causal states does a better job, i.e. gives more information. This property can be formulated as a mathematical theorem which is called the maximal predictability theorem. This is stated as if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
$H[{\overrightarrow{\mathcal{S}}}^{L}|R]\geq H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}].$
Where, $H(\cdot|\cdot)$ is the conditional entropy, $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future. Thus, the inequality shows that the uncertainty of $\overrightarrow{\mathcal{S}}^{L}$ given the causal state $\mathcal{S}$ is less than all other states.
After achieving optimal predictiveness, the causal state set remains the one with minimal statistical complexity. We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$ which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
Next, we present the minimum statistical complexity theorem: for all prescient rivals $\hat{\mathcal{R}}$,
$C_{\mu}(\hat{\mathcal{R}})\geq C_{\mu}(\mathcal{S})$ (2)
We next show that the causal states are minimally stochastic. That is to say, compared with other competitors with the same ability of predicting the future, the causal states and their transition dynamics, have the least uncertainty. Then we have minimal stochasticity theorem which is expressed as: for all prescient rivals $\hat{\mathcal{R}}$,
$H[\hat{\mathcal{R}^{\prime}}|\hat{\mathcal{R}}]\geq H[\mathcal{S}^{\prime}|{\mathcal{S}}].$
Where $\mathcal{S}^{\prime}$ and $\hat{\mathcal{R}}^{\prime}$ are the next causal state of the process and the next state, respectively. This means causal state and $\epsilon$-machine gives the best intrinsic determinism.
Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from observed data? The authors introduced a hierarchical machine reconstruction algorithm in Crutchfield (1994), details are ignored.
Although this algorithm may not be applicable in all operational scenarios, the authors have presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
It is interesting to compare the theory of computational mechanics with causal emergence. Indeed, we can understand all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states. Thus the function $\eta$ that can map a history $\overleftarrow{s}$ to a state $R$ is a possible coarse-graining strategy.
It is worth to be pointed out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can has at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history. Therefore, $\epsilon$ is similar to the notion of effective coarse-graining strategy in reference  Zhang and Liu (2022) (see section 4.1) and the causal transitions $T$ are the corresponding effective macro-level dynamics. And the feature of minimal stochasticity characterize the deterministic property of the macro-dynamic. This is characterized by effective information (EI) in causal emergence theory.
Although the clear definition and quantitative theory of emergence are not given, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003). In Crutchfield (1994), the authors explain that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer. Additionally, they differentiate intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.
3.3 G-emergence Theory
G-emergence theory proposed by Seth in 2008 Seth (2008) is one of the earliest works on quantitative measure of emergence. His basic idea is to use nonlinear granger causality to measure weak emergence in complex systems.
Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$ along, then $B$ is said to be "Granger-cause" of $A$. This implies the existence of Granger causality between $A$ and $B$.
When applying a bivariate autoregressive model for predictions, residual terms are included in the equations of the two variables. Then, the residuals can be utilized to quantify the extent of the causal effect in G-causality. The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals, one is residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other one is the residual of the full prediction model, which shown in Figure 3. Besides, in Seth (2008), the author also defined G-autonomous as if the past values of a time series can help to predict its own future values. And the degree of G-autonomous can be measured in a similar way as G-causality.
Figure 3: G-causality and G-autonomous
With these two basic notions in G-causality, the author defines G-emergence on macro-variables in Figure 4: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if: (i) $M$ is G-autonomous with respect to $\{m\}$ and (ii) $\{m\}$ is G-cause of $M$. The degree of G-emergence can also be quantitatively measured by multiplying the degree of G-autonomous of $M$ and the average of G-causes of $\{m\}$ on $M$.
Figure 4: G-autonomous
The author tests G-emergence theory on the Boid model. This model is a famous artificial life model to simulate the flocking behavior of birds by using three simple rules: cohesion, alignment, and separation Reynolds (1987). All these basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled. The author found that with the strength of cohesion increased, the G-emergence of the entire flock is also increased. Here, the macro-variable is selected as the center of the flock (center-of-mass), and each bird is a micro-variable. The author also discovered a downward causality phenomenon in this simple model that is the center of mass can be used to predict each individual bird. However, the author does not distinguish downward causation from other common causality in this work because Granger’s causality is not a real causal relationship.
Seth’s G-emergence theory is the first attempt to quantify emergence phenomenon via causality measure. However, the causality measure that the author used is granger causality which is not a strict causal measure and it is also depends on the regression method to be used. Furthermore, the measure is defined on variables but not dynamics which means the result depends on the selection of variables.
3.4 Other Quantitative Theories of Emergence
There are alternative means of measuring emergence that do not require relying on causality. Two different ways are discussed by the scholars, one is to understand emergence as a process from disorder to order, another one is to understand emergence from the view of "the whole is greater than the sum of parts".
For example, Moez Mnif and Christian Müller-Schloer used Shannon’s entropy  Mnif and Müller-Schloer (2011) to measure order and disorder. In a self-organized process, emergence occurs when there is an increase in order. This increase can be quantified by measuring the Shannon entropy difference between the initial state and the final state, denoted as $H_{start}-H_{end}$. However, this definition has two limitations. Firstly, the measurement of entropy, denoted as $H$, is dependent on the abstract level of observation. Therefore, it is necessary to account for the entropy increase resulting from a change in the observer’s abstract level. Secondly, the choice of the initial condition of the system is arbitrary. To address this limitation, one approach is to measure the relative level of Shannon entropy compared to the maximum entropy distribution. Finally, emergence can be quantified by:
$Emergence=H_{max}-H-\Delta H_{view},$ (3)
where, $H_{max}$ is the maximized entropy of the system. If no prior information is available, the maximum entropy, denoted as $H_{max}$, corresponds to the Shannon entropy of an equal probability distribution. On the other hand, $H$ represents the Shannon entropy of the system at the final moment of a self-organization process. Additionally, $\Delta H_{view}$ represents the entropy increase during this process resulting from a change in the observer’s abstraction level. It should be noted that if the observer does not alter their abstraction level, this term would be zero. We can also normalize this quantity by dividing $H_{max}$ such that different features can be compared by the normalized quality. For a multivariate system, the authors suggest employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables. Then the authors apply this method to a simulated system of chickens. M. Tang and X. Mao applied this indicator on artificial society models  Tang and Mao (2014).
Inspired by Moez Mnif and Christian Müller-Schloer’s work,  Fisch et al. (2010) suggests to use the divergence measure between two probability distributions to better quantify emergence. They understand emergence as an unexpected or unpredictable change of the distribution underlying the observed samples. However, this method has the defect of computational complexity and estimation accuracy. To conquer the problems,  Fisch et al. (2012) further proposed an approximating method by using Gaussian mixture model to estimate the density and introducing Mahalanobis distance to characterize the divergence between data and Gausian component, better results can be obtained. In Fisch et al. (2011), the authors systematically compare the three mentioned methods, and apply them on a simple test example. Another Shannon entropy based emergence measure is given by Holzer and de Meer et al. Holzer et al. (2008); Holzer and de Meer (2011). They consider a complex system as a self-organization process, in which different individuals interact with communications. Then emergence can be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
Different from all the mentioned methods above, Teo et al. (2013); Szabo and Teo (2015) proposes a method to quantify emergence based on the idea of "the whole is greater than the sum of its parts", and it defines emergence from the interaction rules and states of agents instead of on the overall statistical measure of the whole system. Specifically, this measure consists of two terms that are subtracted from each other. The first term characterizes the collective states of the entire system, whereas the second term represents the summation of the individual states of all its constituent parts. This measure emphasize the emergence that arises from the interactions and collective behavior of the system. Then this method is tested on the example of bird flocks simulation.
3.5 Erik Hoel’s Causal Emergence Theory
3.5.1 Basic idea
The first quantitative emergence theory defined on Markov dynamics and causality measure based on intervention is Erik Hoel’s causal emergence theory.
In this framework, system properties can be characterized at various levels, ranging from micro to macro. If a system exhibits stronger causality at the macro level than at the micro level, it demonstrates causal emergence. Causality is reflected between successive states during the system’s evolution. The strength of a system’s causality reveals the extent to which its future state is influenced by its current state. The basic idea of causal emergence can be illustrated by Figure 5.
Figure 5: The basic idea of Erik Hoel’s causal emergence theory.
For example, statistical mechanics is a typical theory of causal emergence. On the micro-level, a huge number of molecules collide each other and exhibit random behaviors such that probabilistic language must be used to describe them. However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure and temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors. Therefore, thermodynamic laws have stronger causal effect than micro-level molecular dynamics.
Formal tools are used to describe the elements in Hoel’s causal emergence theory. Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro systems, and the inherent strength of causal effect of the Markov model can be measured with effective information (EI) which indicates how effectively a particular state influences the future state of a system.
EI is an intrinsic property of a system’s dynamics which can be quantified by the transitional probability matrix (TPM). A coarse-graining strategy is a function that can map a set of micro-states into a particular macro-state, and therefore a new dynamical model which can be described by TPM can naturally be derived from the micro-level TPM, and the effective information of the coarse-grained model can also be computed. The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase, and at a certain macroscopic scale of coarse-graining, the effective information reaches a maximum: this is the scale at which the system state has maximum causal power to predict future states in the most reliable and effective way.
To illustrate the causal emergence, let’s look at a particular Markov chain model with 4 possible states, whose transition probability matrix (TPM) is:
$S_{m}=\begin{bmatrix}1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 0&0&0&1\end{bmatrix}$ (4)
In this model, if we take the first 3 states as a group 1 and the last state as group 2, anyone in group 1 can transit to all 3 states of the same group with equal probability at the next moment, while because there’s only 1 state in group 2, so the 4th state will always stay in its position. Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty is mainly from group 1.
However, if we merge the first 3 states in group 1 into 1 new state, and keep the 4th state in group 2 as it is, then we get a coarse-grained Markov model with 2 states, with TPM:
[…]
Now the future states of the new system can be fully determined by the current states. It shows that we can eliminate the uncertainty of an undeterministic system by doing coarse-graining over the system states.
3.5.2 About Coarse-graining
The example above illustrates how a system’s determinism can increase as it is coarse-grained. Coarse-graining is a process that simplifies the description of a system by grouping its components into larger, more slowly varying units. It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of microscale interactions. Unlike dimension reduction techniques, such as PCA and SVD, coarse-graining takes into account the system’s features at different spatial and temporal scales. Coarse-graining also differs from renormalization. In physics, the method of renormalization group is invented to avoid the infinity in integrals, it is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro- levels Christensen and Moloney (2005); McComb (2003). However, coarse-graining methohd does not have this requirement in general. Although both techniques are targeting to describe a system from a coarse-grained level, however, coarse-graining are more focused on the states of the system, while renormalization cares of the dynamics, the system rules or partition function etc, which is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking. These are aspects that coarse-graining typically does not consider.
Given a transitional probability matrix, there are a large amount of ways to partition the state space in order to coarse-grain the system. What kind of coarse-graining strategies are more reasonable? In Hoel et al. (2013) and Hoel (2017), the authors suggest to select an appropriate coarse-graining method by maximization effective information of the coarsed TPM. In history, there is another criterion to select coarse-graining strategy called lumpability. Where lumpability of a TPM refers to the coarse-grained TPM exhibits similar dynamics and observation statistics Kemeny and Snell (1983). Commutativity is another requirement which needs the operations of coarse-graining and state transition being commutative Pfante et al. (2014). There are also other discussions or reasonable model reduction of Markov or hidden Markov models Kotsalis et al. (2008); White et al. (2000); Wolpert et al. (2014)
3.5.3 About Effective Information
In order to quantify emergence, a widely used method is to calculate the mutual information between the past states and future states of a system. This mutual information defines the upper limit of the information being transferred from past to future. When the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
$I(X_{t};X_{t+1)}=\sum P(X_{t},X_{t+1})\log\frac{P(X_{t},X_{t+1})}{P(X_{t})P(X_{t+1})}$ (6)
One limitation of utilizing mutual information is that the resulting value may fluctuate in response to changes in the joint probability of data. This can make it difficult to obtain a consistent outcome when the system is exposed to different inputs.
To address this limitation, Hoel introduced the concept of effective information. This concept measures the causal effect of the current state of a system on its future state, also based on mutual information. Effective information is a scoring metric that is calculated from the system’s transition probability matrix (TPM). The TPM does not change solely based on input data.
Assuming that the current state of the system is in the maximum entropy distribution $I_{D}$ (uniform distribution), the TPM predicts that the state distribution of the system at the next moment (i.e., in the future) will be $E_{D}$. The effective information $EI$ of the system is then defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
$EI=I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{X_{t}\sim U}$ (7)
$X_{t}\sim U$ represents the probability of the random variable $X_{t}$ following a uniform distribution. It can be readily demonstrated that the effective information ($EI$) can be fully described by the transition probability matrix ($TPM$) (see Appendix 7).
[…]
Effective information offers a solution to the aforementioned limitation of mutual information. Given that TPM can capture the inherent nature of a system, effective information is likewise an intrinsic property of the system.
Causal emergence involves assuming a uniform distribution for the input data, which can be seen as a way of manipulating the input data to isolate the interference of the original distribution of input data. This manipulation is formalized using Judea Pearl’s causation theory, specifically through the "do" operation. By imposing this assumption of a uniform distribution, we create a controlled environment with equal probabilities of $X_{t}$’s states and allows us to better understand the causal relationships between $X_{t}$ and $X_{t+1}$. Therefore, the term "causal emergence" is used.
We can calculate EIs for both TPMs of macro- and micro- dynamics, and their difference is defined as the measure of causal emergence, that is:
$CE=EI(TPM_{M})-EI(TPM_{m}),$ (9)
where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the one of micro-level. $CE$ measures the degree of causal emergence. If $CE>0$, then causal emergence occurs, otherwise, does not occur.
However, there is a limitation for EI and CE when comparing two dynamics that differ significantly in size. This is because the value of effective information relies on the number of possible states within the system, denoted as $N$, with an upper bound of $\log_{2}(N)$. To facilitate the comparisons between different coarse-graining strategies and scales, effective information is often normalized resulting in a metric known as the effect coefficient $Eff$.
$Eff=\frac{EI}{\log_{2}(N)}$ (10)
The value of effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states. If the information is fully transferred, the effect coefficient is $1$.
In addition to characterizing causal emergence, effect coefficient can be further broken down into two meaningful components: "Determinism," which represents the certainty of a current state evolving into a certain state, or diverging into multiple states in the next moment, and "Degeneracy," which represents the possibility of multiple current states converging into one state in the next moment.
$Eff=Determinism-Degeneracy.$ (11)
It should be noted that the determinism of a system is always greater than its degeneracy, as the lower bound of the effect coefficient is 00. The following examples illustrate what determinism and degeneracy look like in systems with varying TPMs.
Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graning can change them. (a) Deterministic and zero degenerative system; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from c or d, resulting in a fully deterministic system with high effect coefficient.
In Figure 6, the square cells represent for the elements of TPM matrix, and the grey scale represents for the value of TPM elements. Example (a) is a bijective system, meaning that all information from current states can be transferred to future states without loss. It is fully deterministic with zero degeneracy. Example (b) is an extreme case where all current states lead to only one future state, illustrating that high determinism does not necessarily imply a high effect coefficient. Systems (c) and (d) differ but have the same effect coefficients. Finally, system (e) is a coarse-grained version of either system (c) or (d), demonstrating two important points: first, different microscopic systems can be coarse-grained to the same macroscopic system, and second, causal emergence can be quantitatively captured from coarse-graining by the increased effect coefficient.
Despite the fact that effective information can identify and quantify causal emergence, as previously mentioned, there are other independently developed measures of causation, such as Humean constant conjunction Hume (2016), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009) etc. Hoel found all these measures, including effective information can be related to a small set of "causal primatives" Comolatti and Hoel (2022): the probability representations of sufficiency and necessity (and its variant) of cause $c$ and effect $e$.
$suff(e,c)=P(e|c)$ (12)
$nec^{\dagger}(e,c)=P(e|C)$ (13)
$-\𝑐nec(e,c)=1-P(e|C\verb|\|c)$ (14)
Where $C$ stands for the set of all causes and $-\𝑐C\verb|\|c$ stands for the set of all causes but $c$. Hoel pointed out that determinism is actually a measure of sufficiency, while degeneracy is an inverse measure of necessity. The consistency in causal primitives indicates there is a scientific foundation of these individually developed measures of causation.
One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation 8. Therefore, Varley (2022) proposed the conception of flickering emergence to decompose EI by each term being added in Equation 8. This new feature can characterize the local properties of a Markov dynamics.
3.5.4 Examples of Causal Emergence
The example above demonstrates a coarse-graining strategy that is intuitive, as it involves aggregation at the state level. However, in reality, coarse-graining often occurs among variables, and the combination of these variables produces different states. Let’s look at a more complicated example of a boolean network that given in Hoel et al. (2013) and we will inspect the causal emergence when coarse-graining its state transition mechanism. A, B, C and D are 4 binary variables whose state transational relations are as shown in Figure  7(a).
Figure 7: The mechanism and transition probability matrix of a boolean network. Each variable’s next state is determined by the combination of the current states of two other variables, as shown in the boolean network (a). If we examine an atomic subsystem with two inputs and one output, such as AB -> C, the next state of C depends on the present states of A and B. If A=0 and B=0, then the probability of C being 1 is 70%, while the probability of it being 0 is 30%, as indicated in table (b). This same microscopic dynamics applies to other atomic subsystems such as AB -> D, CD -> A, and CD -> B. When we expand to a subsystem of two inputs and two outputs, AB -> CD and CD -> AB, we obtain the transition probability matrix (TPM) demonstrated in table (c). Finally, we extend to the dynamics of the entire system ABCD -> ABCD, whose TPM is represented in a heatmap (d). The effective information and relative metrics of the entire system are as below: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
[…]
Figure 8: An example to show how different coarse-graining strategies on the same boolean network shown in Figure 7 can lead to different boolean dynamics. (a) Coarse-graining strategy 1. (b) Macro-dynamics leading by strategy 1. (c) The corresponding heatmap of TPM and effective information metrics of the macroscopic system by method 1. (d) Coarse-graining strategy 2. (e) and (f) are the macro-dynamics and TPM correspond to the coarse-graining strategy 2 in (d) respectively.
The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.05$, respectively. It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), while it has decreased in the second case ($EI_{ma2}(0.05)<EI_{mi}(1.15)$). Thus, causal emergence is observed in the first coarse-graining instance. This implies that 1) proper coarse-graining can lead to causal emergence in a macroscopic system, and 2) not all coarse-graining approaches can result in higher effective information. This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well defined quantification method. This gives us the opportunity to identify causal emergence in a given system. There is a question coming next: what coarse-graining strategy can generate the maximum effective information? We will address this question in section 4.
3.6 Extensions and Applications
We will discuss the extensions and applications of Hoel’s theory of causal emergence in this sub-section.
3.6.1 Extensions on Continuous Systems
Effective information and its associated metrics have been shown to be useful in quantifying causal emergence. However, there are some limitations to be aware of. First, the method is only applicable to discrete-state systems. Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate effective information metrics. Efforts have been made to apply or extend the concepts of effective information to continuous systems, such as Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
Ordinal Partition Network
One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones by ordinal partition network (OPN) method Varley (2020). The basic idea is to transform the time series data generated by a continuous dynamical system into sub-time series of dimensionality $d$ with discrete values. And then to establish a Transition Probability Matrix (TPM) between these discrete series. This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series. By employing rank orders, this method ensures that the values are restricted to integers within a certain range.
In Varley (2020), the author applied OPN method on Rossler chaotic attractor system. By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) is sensitive to the critical phase transition between period and chaos Varley (2020).
Causal Geometry
Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the conception of causal geometry Chvykov and Hoel (2020).
To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is:
$y=f(x)+\varepsilon,$ (15)
where $x$ is the input variable defined on the interval $[-L/2,L/2]$ and $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a gaussion noise with zero mean and standard deviation $\epsilon\ll L$. Then, an approximate form of the effective information for such continuous system can be obtained by Gaussian integral.
$EI\approx\ln(\frac{L}{\sqrt{2\pi e}})+\frac{1}{2L}\int_{-L/2}^{L/2}\ln\left(\frac{f^{\prime}(x)}{\epsilon}\right)^{2}dx.$ (16)
The reason why we set the domain of $x$ on $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented, and also facilitate to define a uniform distribution on it.
This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ as $x(t)$ and $x(t+1)$ respectively. The formula also can be extended to higher dimension. Suppose $\mathbf{x}\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers, Equation 16 can be generalized to the following form:
$EI\approx\ln(\frac{L^{n}}{(2\pi e)^{m/2}})+\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2},$ (17)
where, $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, and ||$\cdot$|| ⋅ | is the operation to take absolute value, and $\det$ is to take the determinant.
Notice that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$, thus the term in the expectation in Equation 17 can be written as:
$\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2}=\left|\det\left(\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right]\right)\right|,$ (18)
and this is the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
$g_{\mu\nu}\equiv-\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right],$ (19)
it measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ on $\mathbf{x}$. Where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$. Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, and it is the collections of all possible distributions of $p(\mathbf{y}|\mathbf{x})$. This is where the term of “geometry” in this framework comes from.
Finally, we can obtain an expression for EI with Fisher information metric:
$EI\approx\ln\frac{L^{n}}{(2\pi e)^{m/2}}-\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln|\det(g_{\mu\nu})|.$ (20)
This formula may be generalized to the cases where $p(\mathbf{y}|\mathbf{x})$ is non-Gaussian distribution. Once the distribution function is known, we can obtain its Fisher information metric, then EI can be calculated via Equation 20. The reason behind is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
While, in Chvykov and Hoel (2020), the authors consider a more general case, in which the intervention noise is added on the input(intervention) variable $x$ in Equation 16. The intervention noise is $\xi\sim\mathcal{N}(0,\delta^{2})$, and $\delta$ is the standard deviation of $\xi$. While, $\varepsilon$ is called observational noise to differentiate with $\xi$, and it is added on the output(observational) variable $y$ according to Equation 16. Finally, if two kinds of noises are both considered, and if $L=1$ and $\epsilon\ll 1$, Equation 16 becomes:
[…]
This is the formula of EI for continuous mapping function (Equation 15) given by Chvykov and Hoel (2020). Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation 15) and a more discrete map with an adjustable parameter in the logistic function. They discovered that when the noise level is low, the continuous map may exhibit higher EI than the discrete map. However, as the noise level increases, discretizing the mapping function can lead to a model with higher EI. This phenomenon helps to explain why digital circuits eventually outperform analog circuits in mitigating noise interference: the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we can not control $\mathbf{y}$ by intervening $\mathbf{x}$ directly. Instead, we can intervene $\mathbf{x}$ to influence $\theta$, and to affect $\mathbf{y}$ indirectly. Therefore, the three variables form a Markov chain: $\mathbf{x}\rightarrow\theta\rightarrow\mathbf{y}$.
[…]
Finally, the EI calculation formula for the causal geometry is:
$EI_{g}=\ln\frac{V_{I}}{(2\pi e)^{n/2}}-\frac{1}{2V_{I}}\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}\ln\left|\det\left(I_{n}+\frac{h_{\mu\nu}}{g_{\mu\nu}}\right)\right|d^{l}\theta,$ (22)
in which, we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
EI calculation for neural networks
One of the application areas for continuous EI is on neural networks. In Marrow et al. (2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compare the results with information bottleneck theory. The method to calculate EI that they used is to convert continuous variables into discrete ones by dividing the domains of input and output variables into small regions. However, this method has high computational complexity, and is hard to extend to high dimensions.
In Zhang and Liu (2022), the authors propose a new numeric method for calculating EI on neural networks. The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping following:
$\mathbf{y}=f(\mathbf{x})+\varepsilon,$ (23)
where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, and $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network, $\varepsilon$ is a Gaussian noise with zero means and the covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ on the whole training data set. In this way, the neural network can be regarded as a Gaussian distribution $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
Therefore, Equation 17 can be used on this neural network to calculate EI. In Zhang and Liu (2022), the authors use Monte Carlo integration method to estimate the expectation in Equation 17, this technique reduces the computational complexity in a large degree.
However, when the authors compared the results calculated by Equation 17 for different neural networks with different dimensions, they obtained unreasonable results: EI increases with the number of dimensions, thus, causal emergence is always not be observed because micro-dynamics has unreasonably larger EI. Another defect of Equation 17 is that the parameter $L$ always dominates the value of the expression which should be removed from the equation. One way to solve this issue is to use Eff instead of EI (see Equation 10), however, not all $L$s in the Equation 17 can be eliminated. Therefore, the authors invented a notion of dimension averaged EI,
$dEI\equiv=\frac{EI}{m}$ (24)
which simply divide EI by the dimension of the output variable $\mathbf{y}$. By using dimension averaged EI, $L$ can be also removed if we calculate dimensional averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ for macro- and micro- dynamics respectively, that is:
$dCE(f_{M},f_{m})\equiv\frac{EI(f_{M})}{n_{M}}-\frac{EI(f_{m})}{n_{m}},$ (25)
where, dCE is the dimension averaged causal emergence measure. If we bring Equation 17 into Equation 25, we get:
$dCE\approx\frac{\mathbb{E}_{\mathbf{x}_{M}}\log|\det(\partial_{\mathbf{x}_{M}}f_{M})|+\sum_{i=1}^{x_{M}}\log\sigma_{i,M}^{2}}{n_{M}}-\frac{\mathbb{E}_{\mathbf{x}_{m}}\log|\det(\partial_{\mathbf{x}_{m}}f_{m})|+\sum_{i=1}^{x_{m}}\log\sigma_{i,m}^{2}}{n_{m}},$ (26)
in which, $L$ is eliminated.
3.6.2 Causal emergence in complex network
Many complex systems can be represented by networks, therefore, to apply the framework of causal emergence on networks is necessary. However, two problems should be solved in advance to apply Hoel’s framework, one is to assign a dynamics on the studied network because networks are static structures without any dynamical properties, another one is to coarse-grain the network.
Klein and Hoel Klein and Hoel (2020) solved the first problem by considering random walks on complex networks. Subsequently, the TPM of the system is defined based on the transfer probabilities between nodes of a large number of random walkers. Due to the good mathematical properties of the random walk dynamics on graphs, the authors can establish an explicit expression of EI on networks, and the final form only relies on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network. The expression is as follow:
$EI=H(\langle W^{out}_{i}\rangle)-\langle H(W^{out}_{i})\rangle.$ (27)
Where, $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,...,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes which characterizes the degeneracy of the dynamics. And $H(W^{out}_{i})$ is the uncertainty of node $i$.
To tackle the second problem, the greedy algorithm is used to group nodes to form a macro network in Klein and Hoel (2020). It is worth noting that when merging micro nodes into macro nodes, the TPM of the corresponding macro network can be derived by merging the probabilities in TPM of the micro-level. To ensure that the grouped macro network maintains the same random walk dynamics as the original one, dynamic consistency test is implemented.
In the experimental section, the paper explores the application of greedy search for identifying causal emergence in complex networks. The experiments are conducted on artificial networks, as well as four types of real networks. In the case of ER random networks, the size of effective information is solely dependent on the connection probability $p$, and converges to $-\log_{2}p$ with the increase of network size. Additionally, a significant finding reveals the presence of a phase transition point, where the average degree of the network reaches approximately $\log_{2}N$. Beyond this point, the random network structure does not contain additional information with increasing size. In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increases as the network size expands. Conversely, when $\alpha>1.0$, the opposite conclusion is reached. The scale-free network corresponding to $\alpha=1.0$ represents the critical boundary of growth. Regarding real networks, the authors find that biological networks exhibit the lowest EI due to the presence of significant noise, which can be removed by effective coarse-grained, and causal emergence is the most significant compared with other types of networks. However, technical networks exhibit sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships. Consequently, they possess the highest EI among the networks studied.
The network coarse-graining method mentioned above is the greedy algorithm. However, when the network is very large, the efficiency of this method remains considerably low. Following that, Griebenow et al. Griebenow et al. (2019) introduced a spectral clustering-based approach to identify causal emergence within networks. More specifically, the method involved performing eigenvalue decomposition of the TPM, followed by constructing a similarity matrix using the eigenvectors of the nodes. The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro node. Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyper-parameter $\epsilon$. In this paper, the author additionally proposed a gradient descent algorithm based on deep learning. This approach encompasses several steps, beginning with the random initialization of a grouping matrix. Then this matrix is used to construct a macro network by combining the micro network. Finally, the grouping strategies are automatically learned through the process of maximizing effective information within the macro network. However, this method often falls into the local optimal solution.
3.6.3 Other Applications
Once the method for quantifying causal emergence in complex systems is developed, it can be applied across various fields that possess abundant network data. The first type of networks which was studied is biological networks.
As previously discussed, biological networks are full of noise, which poses challenges in comprehending their internal operating principles. On the one hand, such noise arises from the inherent fluctuations within the system itself, while on the other hand, it can be introduced through measurement or observation processes. Consequently, Klein et al. Klein et al. (2022) further explored the relationship among noise, degeneracy and certainty in biological networks and their specific meanings. For instance, in gene expression networks, highly deterministic relationships indicate that the expression of one gene almost invariably leads to the expression of another gene. Simultaneously, degeneracy is a prevalent phenomenon in the evolutionary processes of biological systems. Due to these two factors, it remains unclear at which scale biological systems should be analyzed to gain a deeper understanding of their functions.
To address this, Klein et al. Klein et al. (2021) conducted an analysis of protein interaction networks across more than 1,800 species. They employed EI as a measure to assess the levels of noise and uncertainty in protein interactions. The findings revealed that the macro-scale network exhibited lower levels of noise and degeneracy. Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to nodes that were not part of the macro-scale interaction group. Through robust analysis, the authors demonstrate that eukaryotes exhibit a stronger degree of causal emergence compared to archaea. Additionally, to address the ’deterministic paradox,’ the authors introduce the concepts of the neutral process and the selection process in biological evolution. The neutral process operates at the micro scale, leveraging mutations to promote interactions and enhance species diversity. On the other hand, the selection process operates at the macro scale, effectively eliminating noise that hampers system operation and efficiency. Therefore, in order to adapt to the demands of evolution, it becomes essential for evolved biological systems to function across multiple scales.
Hoel et al. Hoel and Levin (2020) conducted further research on the causal emergence within biological systems. The authors elaborated that macro and micro systems exist widely in biological systems, for example, the micro scale of a group of cells can be the potential ion channel changes, and the macro scale corresponds to the membrane potential changes of cells. Furthermore, the authors utilized EI to analyze gene regulatory networks, aiming to identify the most informative models for controlling mammalian heart development. By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it is revealed that informative macro-scale structures are prevalent across biological systems. Additionally, the authors emphasize the importance of evolutionary systems operating at multiple scales, as it offers significant advantages. Natural selection requires variation between populations, and degradation observed in biological systems serves as a crucial factor for evolution. Degradation provides the conditions for the evolutionary process. However, organisms also need to maintain predicted consistency in phenotype, behavior, and structure to ensure survival and reproduction. Consequently, evolved systems need to operate at multiple scales, and the function of multi-scale systems is malleable in changing environments.
Swain et al. Swain et al. (2022) conducted an investigation into the impact of ant colony interaction history on task assignment and task switching. They employed effective information to examine how noise information propagates among ants, while also exploring the relationship between EI and the proportion of ants assigned to different tasks within ant colonies. The study revealed that the extent of historical interaction between ant colonies influenced task assignment. Additionally, the specific type of ant colony involved in an interaction determined the level of noise present within that interaction. For example, the interactions between foragers displayed significantly higher levels of noise when contrasted with interactions between nurses or cleaners. Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also play different roles in maintaining group cohesion.
EI indicator and causal emergence theoretic framework can also be applied to artificial systems. For example, Marrow et al. Marrow et al. (2020) quantify and monitor the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer. By observing the changes of EI, including determinism(sensitivity) and degeneracy, throughout model training, the generalization ability of the model can be determined, thus helping us to better understand and explain the working principle of neural networks.
Varley et al. Varley (2020) attempted to apply the causal emergence framework to both discrete cellular automata and continuous Rossler systems. In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic and complex. By considering each state as a node, where each state determines the subsequent state, a directed state transition graph is constructed. The analysis revealed that rules 1, 2, and 4 correspond to the strongest causal emergence of dynamics. Notably, the network constructed by these three rules exhibited a significant presence of star and spoke motifs. In addition, the authors also draw some quantitative conclusions, for example, among the 17 rules belonging to the third and fourth categories, 30% exist causal emergence, 70% show causal degradation, and the CE of cellular automata with the same rules remained relatively consistent across different sizes.
Furthermore, the author employed the OPN algorithm to transform the continuous system into a discrete state transition graph for comparative analysis (see section 3.6.1). They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
3.7 Critiques
In the extensive literature on causality and emergence, Hoel’s theory has gained attention for its association of emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner. However, Dewhurst  Dewhurst (2021) provides some philosophical clarity of Hoel’s theory, arguing that it is epistemological rather than metaphysical. It suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving "genuinely novel causal powers." This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
The computation of Hoel’s effective information relies on two premises: 1) knowledge of the system’s microscopic dynamics, and 2) knowledge of the coarse-graining scheme. However, in practice, it is rare for both conditions to be simultaneously satisfied, especially in observational studies where both are unknown. This limitation hinders the practical applicability of Hoel’s theory.
It is pointed out that Hoel’s theory negleted the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022). Additionally, the combination of some coarse-graining operations over states and coarse-graining operations over time does not exhibit commutativity. For instance, assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold. This indicates that certain coarse-graining operations can result in a discrepancy between the evolution of macroscopic states and the coarse-grained states of the evolved microscopic systems. It implies the need for consistent constraints on coarse-graining strategies.
That means solely maximizing EI may raise some problems, and further constraints must be added in the framework. We will discuss this problem in the section of 4.1.2.
3.8 Rosas Fernando’s Quantifying Causal Emergence
3.8.1 Basic idea
In Hoel’s framework, it is essential to find a coarse-graining strategy in order to determine the occurrence of causal emergence, and the outcome is influenced by the choice of the coarse-graining method. While it has been suggested by Hoel Hoel et al. (2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised as proposed by Dewhurst Dewhurst (2021). Third, EI is an global measure because it requires to intervene the input variable $X$ to be a uniform distribution on the whole domain Rosas et al. (2020). However, many regions are not be able to be observed from data. Consequently, there is a pressing need for an alternative theoretical framework of causal emergence that does not rely on a coarse-graining method.
In response, Fernando E. Rosas took an approach that did not require a coarse-graining strategy as prerequisite, attempting to break down excess entropy - the mutual information between a system’s past and future states - into non-overlapping parts to identify the information components most relevant to causal emergence. To accomplish this, he relied on the Partial Information Decomposition (PID) framework proposed by Williams and Beer, which provides a method for non-overlapping decomposition of joint mutual information Williams and Beer (2010). First, we will introduce Williams and Beer’s theoretic framework.
3.8.2 Partial Information Decomposition
The Partial Information Decomposition (PID) framework investigates the general informational relationship between source variables and a target variable. To make the problem description simpler without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
Figure 9: Venn diagram of PID. (a) Relationship of mutual information between Y and X1, X2. (b-d) Examples that joint mutual information is solely contributed by redundant, synergistic, and unique information. (e) Redundancy lattice representation of (a).
The mutual information between the target variable and individual source variables,$I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibit a complex relationship and can not be intuitively converted to one another. Nonetheless, it’s an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
$I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)+Un(X_{1};Y|X_{2})+Un(X_{2};Y|X_{1})+Syn(X_{1},X_{2};Y)$ (28)
Specifically:
  • $Red$: Redundant Information refers to the information held by both sources;
  • $Un$: Unique Information refers to the information held by one source but not the other;
  • $Syn$: Synergistic Information is the information held by all sources together, but not any individual one.
If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates that they must be independent of one another.
For a more intuitive understanding, let’s consider a few simple toy examples.
Case 1, $Y=X_{1}=X_{2}$: This is a scenario in which a single source variable can predict the target variable, and the addition of another source variable does not enhance the prediction of the target variable. In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, Fig 9(b).
Case 2, $Y=X_{1}\oplus X_{2},X_{1}\perp{X_{2}}$: In another scenario, neither of the source variables can predict the target variable individually, but together, they can predict the target variable synergistically. In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, Fig 9(c).
Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, Fig 9(d).
For general cases, Williams and Beer presented a method for calculating redundant information defined as $Red_{min}$. This method reflects the concept of redundancy, identifying it as the shared information across all sources, is equivalent to the minimum amount of information contributed by any single source:
$Red_{min}(X_{1},X_{2};Y)\sum_{y\in\mathcal{Y}}P(Y=y)\min_{Xi}I(Y=y;X_{i})$ (29)
Fig 9(e) presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010). In this representation, {12} denotes synergistic information, $\left\{1\right\}$ and $\left\{2\right\}$ denote unique information, and $\left\{1\right\}$$\left\{2\right\}$ denotes redundant information. In the subsequent sections, we will also see that the $\phi$ID framework employs the notation of the redundancy lattice.
Although the PID framework remains compatible with scenarios involving more than two source variables, it should be noted that the corresponding Venn diagrams and redundancy lattices for these scenarios can become substantially more complicated and difficult to decipher, as discussed in Williams and Beer (2010).
3.8.3 Integrated Information Decomposition
The PID framework provides a useful framework for analyzing the non-overlapping information composition in a multivariate system. However, its application to causal analysis of dynamic systems is limited by the fact that it only allows for a single target variable. This limitation prevents the framework from fully capturing the transitions of multiple states across time steps. To address this challenge, Fernando E. Rosas etc have developed Integrated Information Decomposition ($\phi$ID) Mediano et al. (2019), which takes its name from Integrated Information Theory (IIT) Oizumi et al. (2014). This extension of PID provides a more comprehensive means for analyzing dynamic systems.
To introduce Rosas’s framework clearly, we only consider a system with two variables. All the definitions and calculations can be generalized to the systems with more variables.
The objective of the $\phi$ID framework is to decompose excess entropy into non-overlapping information components. In a two-variable Markovian system, the excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states. From the causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, while $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent effects.
Rosas first noted that mathematically, the PID framework is symmetric with respect to source and target variables. There are two viewpoints for analyzing the above Markovian system with PID. One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the "cause" to elements), while the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the "effect" to elements). The redundancy lattices of these perspectives are illustrated in Fig 10 (a) and (b), and respectively, are referred to as "forward PID" and " PID".
Figure 10: (a) Redundancy lattice of "forward PID", and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of "backward PID" and its target variable $X^{1}_{t}X^{2}_{t}$.
Next, Rosas introduced the $\phi$ID framework, which consolidates the forward PID and backward PID into a single framework. In this framework, the one-to-many relationship in PID is expanded to include many-to-many relationships. He built full connections between the elements of redundancy lattices of the forward PID and the backward PID. This approach generates 16 relations between the source and target, as depicted by the colored lines in Fig 11 (a). (Each color corresponds to a specific source element of $X_{t}$.) These relations are referred to as "$\phi$ID atoms" and can be represented as vertices in a lattice, as shown in Fig 11 (b). This more complex lattice is referred to as the double-redundancy lattice because it is a "product" of two redundancy lattices (of forward PID and backward PID).
Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID. (b) Double-redundancy lattice, whose vertices correspond to the edges in (a)
In the double-redundancy lattice, each vertex is defined as a "$\phi$ID atom," denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, let $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is demoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$. $\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system. Consequently, the formal definition of Integrated Information Decomposition ($\phi$ID) is:
$\mathbb{E}=I(\mathbb{X_{t}};\mathbb{X_{t+1}})=\sum_{\alpha,\beta\in\mathcal{A}}{I^{\alpha\rightarrow\beta}_{\partial}}$ (30)
The main advantage of $\phi$ID is that it can provide more fine-grained analysis of complex systems than PID by introducing the temporal dynamic. By decomposing the mutual information of the system at different times, $\phi$ID provides another approach for the quantitative study of Causal Emergence.
3.8.4 Reconciling Emergences
Based on the concept of synergistic information in the PID framework, Rosas proposed a quantitative definition of causal emergence and further classified it using the $\phi$ID framework, which addresses the problem of finding a coarse-graining strategy. The definition includes two aspects: the first is whether the system has the ability to produce causal emergence objectively. Then, on this basis, whether there is causal emergence happening with specific macroscopic features.
In terms of the capacity to generate causal emergence for a system, this definition links causal emergence to synergistic relationships among variables in different times. Therefore, a system $X_{t}$ is said to have capacity for causally emergent features if and only if that:
$Syn(X_{t};X_{t+1})>0$ (31)
This means that causal emergence is regarded as the synergy effect among variables between the previous and subsequent moments in the system with Markovian dynamics.
Then, Rosas further classify causal emergence by the $\phi$ID framework. Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ by the $\phi$ID, here are four information atoms corresponding to synergistic effect, which is regarded as causal emergence, are $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. Since $\phi$ID includes forward and backward PID, he further divides causal emergence into downward causality and causal decoupling according to the different characteristics of those information atoms. The downward causality is defined by $\mathcal{D}(X_{t},X_{t+1})$, which encapsulating the information atoms that manifest a synergistic effect exclusively in the forward PID:
$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (32)
Conversely, the causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
$\mathcal{G}(X_{t},X_{t+1}):={I^{\{12\}\rightarrow\{12\}}_{\partial}}(X_{t},X_{t+1})$ (33)
As the $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that
$Syn(X_{t};X_{t+1})=\mathcal{G}(X_{t},X_{t+1})+\mathcal{D}(X_{t},X_{t+1}).$ (34)
Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) represented in double redundancy lattice.
Rosas also gives out a version of causal emergence quantification depending on a macro-variable, i.e., a coarse-graining strategy. If a system has the capacity to generate causal emergence, then there can be some macroscopic features that exhibit causal emergence. A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ if the complete state $X$ of the system at time $t$ is known with perfect precision, which is equivalent to $V_{t}$ is statistically independent of $X_{t+1}$ when $X_{t}$ is given. Then, for a system described by $X_{t}$, a supervenient feature $V_{t}$ is said to exhibit causal causation if:
$Un(V_{t};X_{t+1}|X_{t})>0.$ (35)
For this definition, the system’s capacity of causal emergence is required, where $Syn(X_{t};X_{t+1})>0$, since $Un(V_{t};X_{t+1}|X_{t})\leq Syn(X_{t};X_{t+1})$ holds for any supervenient feature $V_{t}$. Corresponding to the classification of system’s capacity, the Downward Causation exists of a feature variable $V$ when:
$Un(V_{t};$ $X_{t+1}^{1}|X_{t})>0$ (36)
$or$
$Un(V_{t};$ $X_{t+1}^{2}|X_{t})>0$
and the Causal Decoupling exists when
$Un(V_{t};V_{t+1}|X_{t},X_{t+1})>0,$ (37)
which also needs to be premised on the capacity of the system. Furthermore, $V_{t}$ is said to have pure causal decoupling if $Un(V_{t};X_{t+1}^{\alpha}|X_{t})=0$ and $Un(V_{t};X_{t+1}|X_{t})>0$. If all the emergent features exhibit pure causal decoupling, then the system is said to be perfectly decoupled.
Figure 13: Synergistic causes can be further divided into causal decoupling and downward causation.
Although a rigorous quantitative definition for causal emergence is proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems. Also, there are many calculation methods in the PID framework, which makes the quantitative definition determined by the specific quantitative method when applied. Therefore, to solve those problems, Rosas also relaxed the calculation methods of causal emergence, thus obtaining sufficient conditions for the two phenomena as identification criteria.
Specifically, in order to avoid exploring the specific quantitative relationship of synergistic and redundant information, this criteria repeatedly subtracts redundant information to make positive results a sufficient condition for emergent phenomena. Although this makes the quantitative method lose some of its generalizability, it improves the reliability of the results. The three indicators to be used are:
$\Psi_{t,t+1}(V):={I(V_{t};V_{t+1})}-\sum_{\begin{subarray}{c}j\end{subarray}}I(X_{t}^{j};V_{t+1}),$ (38)
which measures the mutual information between macro-variables across two time steps minus the information between micro-state to macro-state, where $V$ is a predefined macro-variable, but how to find such macro-variable is never discussed in Rosas et al. (2020).
$\Delta_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}\left({I(V_{t};X_{t+1}^{j})}-\sum_{\begin{subarray}{c}i\end{subarray}}I(X_{t}^{i};X_{t+1}^{j}))\right),$ (39)
which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$ in next step minus all the summation of mutual information between $X_{t}^{i}$ and $X_{t+1}^{j}$.
$\Gamma_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}{I(V_{t};X_{t+1}^{j})},$ (40)
this is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$
The specific functions of the indicators are as follows: $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$. Similarly, $\Delta_{t,t+1}(V)>0$ is as sufficient condition for $V_{t}$ to show downward causation. Finally, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ is a sufficient condition for causal decoupling to occur.
In summary, Rosas proposes an approach to quantitatively describe causal emergence by establishing the relationship between causal emergence and the synergistic effect of variables in different times, and further classified causal emergence by $\phi$ID. The definition not only objectively judges whether the system has the capacity to produce causal emergence, but can also measure the characteristics of causal emergence of specific macro features. The most notable contributions of his work were to bring the study of emergence closer to quantitative and empirical scientific research, to classify types of causal emergence, and to complement philosophical discussions about emergence.
There are some applications of PID or $\phi$ID. Recently, Luppi et al. apply $\phi$ID method on brain BOLD signals Luppi et al. (2022) to identify the synergistic core of the brain. They found synergistic information is for the integration of different parts to a whole, and redundancy is for robustness.
With the development of partial information decomposition techniques, we can further decompose the mutual information between two variables. This may help us to reveal the properties of systems with more aspects. For example, in  Varley and Hoel (2022), the authors first decompose the mutual information of a system, and calculate an indicator of synergy bias according to partial information decomposition method Williams and Beer (2010) to characterize how synergistic information distribute on different levels of the system. Higher synergy bias indicates that more partial information is in synergistic relationships. Then, the authors found that in some systems with occurrence of causal emergence, if the system is coarsen to a reduced system, the synergy bias will increase. Therefore, partial information is converted from redundancy to synergy as we coarse-grain the system. The conclusion is emergence is a kind of information conversion.
3.8.5 Comparison with Hoel’s framework
When comparing Hoel’s framework to Rosas’ quantification framework of causal emergence, several clear advantages can be observed in the latter. Firstly, Rosas’ theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal. Secondly, it offers a detailed decomposition of causal emergence, specifically downward causation and causal decoupling. Lastly, it effectively avoids cases of fake causal emergence, where the macro-variable depends solely on unique or redundant information from the micro-variables.
However, there are also some disadvantages to consider. Firstly, in order to obtain the full information lattice, a systematic iteration of all variable compositions is required. Additionally, despite the use of formula 35, it is necessary to define a macro-variable. Unfortunately, the authors do not provide any method to identify such a variable. Secondly, all the mutual information and its decomposition are based on correlations rather than causality. It is crucial to discuss how to incorporate causal elements, such as intervention and counterfactuals, into the framework.
Finally, the issue of identifying whether causal emergence occurs in a system based on the given time series data of its behaviors has not been addressed in the preceding discussion. To tackle this problem, the application of emerging technologies in machine learning and artificial intelligence is required. These technologies can provide valuable tools and techniques for detecting and analyzing causal emergence.
4 Causal Emergence and Artificial Intelligence
Recently, newly emerging artificial intelligence technologies have made significant breakthroughs in tackling a range of important and challenging problems. Examples include defeating human champions in complex games like Go Silver et al. (2016); Dabney et al. (2020), and predicting the intricate structures of protein folding Senior et al. (2020). These achievements have been made possible through the application of machine learning methods, which leverages well-designed neural network architectures and automatic differentiation techniques. Machine learning methods, with their diverse and well-designed neural network structures, along with automatic differentiation techniques, have demonstrated the ability to efficiently approximate complex functions within a vast functional space. Given this capability, it raises the question of whether we can leverage these emerging techniques to address the challenging problem of finding effective coarse-graining strategies or optimal decomposition methods for complex systems using time series data?
The answer is yes. We will introduce two latest works on causal emergence identification from data based on neural networks. Interestingly, these works indicate that causal emergence identification problem can also provide us with insight into connections with another related field, causal representation (machine) learning although causal emergence identification and causal representation learning are totally different research areas for distinct problems from appearance, they share commonalities. Furthermore, we also review the works on causal model abstractions and reinforcement learning based on world models as extensions because similar machine learning architecture are shared by these machine learning frameworks.
4.1 Causal Emergence Identification
In the previous section we have introduced several works on quantifying emergence by causality and other information theoretic conceptions. Among which, all works try to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence. However, in practical applications, in order for the theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time series data of dynamic systems, and give explanations for the results.
4.1.1 Rosas’ Identification Method
The first method for causal emergence identification is given by Rosas’ paper in causal emergence Rosas et al. (2020). Actually, the three indicators, the Equations 38, 39 and 40 as mentioned in section 3.8.4. The identification criteria have been exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
Although these three indicators avoid the problem of redundant information calculation, they are all sufficient conditions, that is, indicators greater than 0 can indicate emergence. The construction of this indicator repeatedly cuts out redundant information, which makes it almost impossible to identify the emergence of this indicator in systems with large redundant information or systems with many variables, and this is the case for most of the systems we face. Another defect of this method is that the coarse-grained variable $V$ must be given in prior, and different $V$ affects the results in a large degree. Therefore, the automatic coarse-graining strategy design according to data is still remain unsolved. New methods are required.
Therefore, the current theoretical frameworks of causal emergence do not give a good workable identification algorithm. Although existing works have tried to propose some workable methods based on static network structures and approximation for information decomposition, a general method for general Markovian dynamic systems is still lacking. One of the important reasons is to identify causal emergence within data needs to search for all the functions of coarse-graining or decomposing subsets for a complex system. However, the conventional numeric methods can not afford the computational cost for such search in a huge functional space. New methods are required.
4.1.2 Neural Information Squeezer (NIS)
Here we introduce one of the latest works on this direction which is just published in this special issue. In this paper, the authors formulate the problem of causal emergence identification on continuous Markov dynamics. Suppose the time series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
$\frac{d\boldsymbol{x}}{dt}=g\left(\boldsymbol{x}(t),\xi\right),$ (41)
Where $\xi$ is a Gaussian noise. The time series $\{\boldsymbol{x}_{t}\}$ are the observations of the micro-states of the system, and the problem that we confront is identifying if the causal emergence occurs for the original system. And we also want to know the appropriate coarse-graining strategies, the macro dynamics which correspond to the coarse-graining method, and the dimension of the phase space on macro-level if the causal emergence occurs.
To solve the problem, the authors propose a mathematical framework by converting the original issue into an optimization problem for functions. That is, we can find the appropriate coarse-graining strategy $\phi_{q}$, which is defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where, $q<p$ is the dimension of the macro-states, and the macro-level Markov dynamics $f$, which is defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$, the dimension averaged EI of the macro-dynamics (see Equation 24), that is,
$\max\mathcal{J}(\hat{f}_{\phi_{q}}),\qquad$ (42)
However, this problem has a trivial solution, e.g. $\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity. Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy within all $q$ dimensional functions. This is trivial because too much information has been abandoned by the coarse-graining strategy such that the macro-dynamic is useless.
To conquer the problem, the authors add a reasonable constraint to the original optimization framework, which is called effectiveness requirement. A coarse-graining strategy and the corresponding macro-dynamic are effective, which means there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state instead of the macro-state in the next time step according to the state in the previous step, the constraint can be written with,
$||\phi^{\dagger}(f(\phi(\boldsymbol{x}_{t})))-\boldsymbol{x}_{t+1}||<\epsilon,$ (43)
for all time step $t$, where, $\epsilon$ is a given constant. In this way, an effective coarse-graining strategy and the macro-dynamic can not only maximize the EI but also reproduce the original micro-dynamic as much as it can.
This constraint is an important complimentary for the framework of EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states which can solve the problem of ambiguity of macro-states. This constraint is also a guarantee of commutativity. To see this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ on both sides of Equation 43 to obtain $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
However, the problem is still hard to solve. The paper Zhang and Liu (2022) proposes a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step, and maximize EI for different $q$ in the second step. The first step can be solved by training a neural network easily, and the second step converts the complex functional optimization problem into a line search in one dimensional space.
Figure 14: The architecture and workflow of the Neural Information Squeezer (NIS). NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$. The encoder $\phi$ comprises an invertible neural network $\psi$ and a projection operator $\chi$, which simply discards fixed dimensions output by $\phi$. On the other hand, the decoder $\phi^{{\dagger}}$ incorporates the same invertible neural network as the encoder but implements its inverse $\psi^{-1}$. In cases where the input is incomplete, it can be complemented using Gaussian random noise.
To realize the abstract mathematical framework, the authors propose an encoder-dynamics learner-decoder framework of neural network, named neural information squeezer (NIS), which describes the basic working principle of the framework vividly. To reduce the complexity of the encoder-decoder which corresponds to $\phi$ and $\phi^{\dagger}$, respectively, the authors use an invertible neural network to implement the encoder, and invert the input-output workflow of the same neural network to realize the computation of the decoder. Concretely, the encoding process is decomposed into two step computation, information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on $p$ dimensional input space, and information discarding with the simple projection operation, that is:
$\phi(\boldsymbol{x})=\chi_{q}(\psi(\boldsymbol{x})),$ (44)
where $\chi_{q}$ is the projection operator to project a $p>q$ dimensional vector into a $q$ dimensional vector with the dimensions from $1$ to $q$ left. The decoding process also contains two stages, random sampling a Gaussian noise to complete the details of macro states, and information conversion but in an inverted way, that is:
$\phi^{{\dagger}}(\boldsymbol{x})=\psi^{-1}(\boldsymbol{x}\bigoplus\zeta),$ (45)
where, $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$ dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.
The adoption of an invertible neural network in the framework serves two main purposes. Firstly, by inverting the workflow of the neural network $\psi$ to obtain the decoder $\phi^{{\dagger}}$, a significant reduction in the number of parameters is achieved. This reduction can lead to improved efficiency and computational benefits. Secondly, the use of invertible functions in the neural network provides favorable mathematical properties that facilitate the analysis of the entire framework in a tractable manner. These properties enable researchers to gain insights and understand the characteristics of the framework more easily.
This architectural design with invertible neural network can not only reduce the complexity of computation, but can also separate the input information from the micro-states into noise and EI for prediction clearly. The authors proved a mathematical theorem stating that, as the neural network is converged by training, the mutual information on the adjacent macro states approaches to the mutual information of the two adjacent-step micro states in the data. This means, all the information dropped by the encoder is almost irrelevant with the prediction. All the useful information is squeezed into the information bottleneck of the dynamics learner at the macro-level. And another theorem has proved that the degree of information squeezing for micro-states is increased as we reduce the dimension of the macro states, and thus more and more useless information in the original micro-state data is thrown out by the narrower and narrower information bottleneck, the macro-state dynamics learner.
To solve the problem of causal emergence identification, we first optimize the neural networks, and the dimension averaged EI, $\mathcal{J}_{q}$ can be obtained for different $q$. When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension averaged EI for the micro-dynamics. Finally, the calculation of
$\Delta\mathcal{J}=\mathcal{J}_{q}-\mathcal{J}_{p}$ (46)
is the measure of dimension averaged causal emergence (see Equation 25. If $\Delta\mathcal{J}>0$, we can judge that causal emergence occurs.
Numeric examples are given to show the effectiveness of the whole framework. Interestingly, the neural information squeezer can work on the time series data with discrete micro-state although the whole framework is built based on continuous dynamical systems. Particularly, the model can work out the same coarse-graining strategy and macro-level dynamic as in the example of boolean network with four nodes in Hoel’s original paper, although any information about the node grouping and state mapping is not given to the framework.
It is also interesting to compare NIS with the mentioned frameworks or models in previous sections. Compare to the theory of computational mechanics, NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as the states, and when the whole framework is well-trained such that it can predict the future micro-states precisely, the encoded macro-states converge to effective states which can be treated as causal states in computational mechanics. However, the objective that maximize the effective information (EI) in NIS has no correspondence.
NIS also has similarity with G-emergence because it also adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step. However, there are several obvious differences between the two frameworks. In theory of G-emergence, the macro-state must be selected by hand at first, however, it is optimized automatically. Second, NIS uses neural networks to predict the future states, while, G-emergence uses auto-regressive techniques to fit the data. Due to the universal approximate theorem(see  Hornik (1991); Teshima et al. (2020)), neural networks are more advantage.
NIS provides a solution for causal emergence identification from data, and it is the first time to solve the problem of the automatic construction of coarse-graining strategy and the macro-dynamics from data for causal emergence identification.
Another important benefit of NIS which is to be noticed is that, what NIS has solved actually is a generalized causal emergence identification problem but not limited on a justification. Because after training, NIS can not only identify if causal emergence occurs, but also can obtain a neural network for coarse-graining the data and a neural network to simulate macro-dynamics. This may provide us richer information than just a justification.
4.1.3 Neural Information Squeezer Plus (NIS+)
One of the biggest problem of NIS is that it does not directly solve the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner). Instead, it searchs for the optimal dimension of the macro-state space $q$ such that EI can be maximized. Therefore, for given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimize Equation 43) but not for EI. Second, NIS can not support complex computations such as the coarse-graining functions containing multiple steps of information conversion and discarding.
To tackle these problems, NIS+ is proposed in ref Yang et al. (2023). Mathematically, the problem of EI maximization can be converted into a machine learning problem according to the definition of EI and the variational inequality and probability reweighting technique. After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},g_{\theta^{\prime}}$ with parameters $\omega,\theta$ and $\theta^{\prime}$, respectively. Formally, the problem of minimization of loss function problem without constraints can be written:
$\min_{\omega,\theta,\theta^{\prime}}\sum_{t=1}^{T-1}w(\boldsymbol{x}_{t})||\boldsymbol{y}_{t}-g_{\theta^{\prime}}(\boldsymbol{y}_{t+1})||+\lambda||\hat{\boldsymbol{x}}_{t+1}-\boldsymbol{x}_{t+1}||,$ (47)
where $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t}))$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t+1}))$ are the macro-states. $\lambda$ is a Lagrangian multiplier which will be taken as a hyper-parameter in experiments. $w(\boldsymbol{x}_{t})$ is the inverse probability weights which is defined as:
$w(\boldsymbol{x}_{t})=\frac{\tilde{p}(\boldsymbol{y}_{t})}{p(\boldsymbol{y}_{t})}=\frac{\tilde{p}(\phi(\boldsymbol{x}_{t}))}{p(\phi(\boldsymbol{x}_{t}))},$ (48)
where $\tilde{p}$ is the new distribution of macro-states $\boldsymbol{y}_{t}$ after intervention for $do(\boldsymbol{y}_{t}\sim U_{q})$, and $p$ is the natural distribution of the data. In practice, $p(\boldsymbol{y}_{t})$ is estimated through kernel density estimation (KDE) Rosenblatt (1956). The approximated distribution, $\tilde{p}(\boldsymbol{y}_{t})$, is assumed to be a uniform distribution, which is characterized by a constant value. Consequently, the weight $w$ is computed as the ratio of these two distributions. Theorems are proved to guarantee the correctness of this conversion. The overall framework of NIS+ can be depicted in Figure 15.
Figure 15: The workflow and architecture of our proposed framework, Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data. (a) Various forms of input data. (b) The framework of our proposed model, NIS+, which incorporates our previous model, NIS. The boxes represent functions or neural networks, and the arrow pointing to a cross represents the operation of information discarding. $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, while $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state. $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})$ represent the macro-states obtained by encoding the micro-states using the encoder. $\hat{\boldsymbol{y}}_{t}=\phi(\hat{\boldsymbol{x}}_{t})$ and $\hat{\boldsymbol{y}}_{t+1}=\phi(\hat{\boldsymbol{x}}_{t+1})$ represent the predicted macro-states obtained by encoding the predictions of micro-states. The mathematical problems that each framework aims to solve are also illustrated in the figure. (c) The various output forms of NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
Additionally, to apply NIS+ to the data generated by the complex systems of multi-agent systems and cellular automata, the structures of encoder can be extended by defining two kinds of combinations: one is to stack the basic encoder, the other is to concatenate the basic encoder. Here, a basic encoder is the composition of an invertible neural network and a projection operation. By inverting the directions of all the basic encoder, and supplements the empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
Numerical experiments demonstrate the effectiveness of NIS+ in capturing emergent behaviors and finding causal emergence in various simulations, including SIR dynamics, bird flocking (Boid model), and the Game of Life. Additionally, NIS+ is applied to find emergent features in real brain data from 830 subjects while watching the same movie. Experiments also demonstrate that the learned dynamical model exhibits improved performance in terms of out-of-distribution generalization compared to other models.
4.2 Implications in Causal Representation Learning and Out-of Distribution Generalization
We found that, as a framework of machine learning, neural information squeezer and plus can be regarded as a kind of framework for causal feature learning Chalupka et al. (2017); Spirtes et al. (2000); Chalupka et al. (2015) or causal representation learning Schölkopf et al. (2021). Causal feature learning aims that to find a high-level causal relations from the micro-level data. In causal representation learning the macro states are causal variables or causal representations Schölkopf et al. (2021), macro-dynamics are causal mechanisms, and the learned coarse-graining strategy is the representation or encoding process. In this sub-section, we will introduce causal representation learning in details at first, and compare the two emerging fields.
4.2.1 Introduction of Causal Representation Learning
Causal representation learning is an emerging field in artificial intelligence, it tries to combine two important fields in machine learning: representation learning and causal inference, together.
Representation learning aims to extract important features (or representations) hidden in data to make predictions automatically Bengio et al. (2013). Deep neural networks are always adopted, important features in data are represented by the activation modes of neurons LeCun et al. (2015); Goodfellow et al. (2016). The modes in different layers usually represent significant features in various scales, and deeper layers always correspond to larger scales Alzubaidi et al. (2021). The features in different layers are combined together to make predictions or classifications. For example, a deep neural network for human face recognition may have two components. The first one is a deep neural network with stacked convolution units LeCun et al. (1998), and it tries to extract features of faces on different scales. The second component is always a common feed-forward neural network which can accept the input features of faces and output a label to identify the character represented by the input image . The first component is for representation, and the second one is for classification. The neural network as a whole must learn to extract important features from images to make high precision classification.
Although representation learning has achieved remarkable success in various domains such as image classification LeCun et al. (2015), face recognition Wang et al. (2018), language understanding Mikolov et al. (2013); Devlin et al. (2018); Vaswani et al. (2017), and game playing Mnih et al. (2015); Dabney et al. (2020), it suffers from a critical limitation: it can only capture the information or associations within the data, but not the underlying causal relationships. According to the theory of causal hierarchy by Judea Pearl Pearl et al. (2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship which is invariant in different environments, while the latter may be more dependent on contexts and reflect the limitation of the data. For instance, in the example of face recognition mentioned previously, if all the images in the training data have outdoor backgrounds, e.g. grasslands, lakes, trees, etc., the neural network may not work well on the images with indoor backgrounds Cui and Athey (2022); Zhang et al. (2021). This is because the neural network may learn the fake association between backgrounds but not foregrounds and labels. However, the association is only the reflection of the bias in the training data, but not the real causal relationship.
In causal inference, people use graphs to represent causal relationships. Nodes represent causal variables, and edges which are always combined with other information like structural equations or conditional probabilities to represent causal relations Pearl et al. (2000). In conventional causal inference, while the parameters in a causal model can be adjusted based on data, all the variables and causal relationships are typically extracted manually rather than directly from data itself.
To conquer the problem, causal representation learning tries to combine the advantages of the two sub-fields to extract the important features and the relationships with causality behind data automatically Schölkopf et al. (2021).
Figure 16: Illustration for the workflow of a typical causal representation learning agent. In the top panel, we have a depiction of a real block world with a robotic arm, along with the corresponding causal graph representing the relationships between the variables in the real world, nevertheless, it is totally unknown for the agent. Additionally, we have the data in the form of images generated by the block world. In the bottom panel, we see the structure of the causal representation learning agent. It utilizes a convolutional neural network (CNN) to extract meaningful representations from the data and perform downward tasks based on these representations. The encoder plays a crucial role in capturing relevant features and patterns from the input data, enabling the agent to learn and understand causal relationships.
A typical scenario of causal representation learning can be depicted by Figure 16. Suppose there are a set of variables and causal mechanisms (including causal graphs and structural equations) between them which can describe how the world of robotic arms and colorful blocks works, whereas, they are not observable for the agent of causal representation learning. What the agent can observe is a set of images which are generated by robotic arms and blocks. The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks.
Formally, if the input data can be described by $X=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{p})$, causal representation learning will construct two functions and a directed acyclic graph, which include a causal representation function: $\tilde{\phi}:\mathcal{R}^{q}\rightarrow\mathcal{R}^{p}$, such that:
$X=\tilde{\phi}(Y),$ (49)
where, $Y=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},...,\boldsymbol{y}_{q})$ is a set of causal variables and $q\ll p$; a directed acyclic causal graph $\mathcal{G}_{S}\equiv(Y,E)$, on which the node set is $Y$, and the directed links in $E$ represent causal relations; as well as a set of causal mechanisms:
$\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i}),$ (50)
for $i=1,2,…,q$, where $f_{i}$ is the structural equation for the $i$th causal variable $\boldsymbol{y}_{i}$, which describes how the parent causal variables $PA_{i}\equiv\{j|j\rightarrow i\in E\}\subset\{\boldsymbol{y}_{i}\}$ and the external unknown variable $U_{i}$ determine the value of $\boldsymbol{y}_{i}$. And there is a requirement for $U_{i}$ for all $i,j$, that is:
$P(U_{i},U_{j})=P(U_{i})P(U_{j}).$ (51)
Thus, any two external unknown variables need to be independent each other.
However, learning the three elements including the two functions and a graph in the same time is very hard, we can convert the problem to reduce the difficulty. This is to learn disentanglement representations, in which the representations are independent(disentangled) each other Wang et al. (2022). We will show how we can derive disentanglement representations from Equations 49 to 51.
Notice that if we combine the equations $\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i})$ for all $i\in[1,q]$ together, we have $Y=F(Y,U)$, where $U=(U_{1},U_{2},...,U_{n})$, and $F=(f_{1},f_{2},...,f_{q})$. That means, the causal mechanisms actually describe how the unknown variables $U$ determine the causal variables $Y$. Suppose we can solve this equation to obtain $Y$ expressed by $U$, and we define $\tilde{F}$ such that: $Y=\tilde{F}(U)$, according to Equation 49,
$X=(\tilde{\phi}\circ\tilde{F})(U)$ (52)
if we define $\phi\equiv\tilde{\phi}\circ\tilde{F}$, then $\phi$ can be understood as a decoder which can decode a disentanglement representation to data. Further, if we could construct a encoding function $\phi^{{\dagger}}:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ which takes $X$ to $U$, and we require that:
$\phi\circ\phi^{{\dagger}}\sim I,$ (53)
where, $I$ is the identity mapping, thus, $\phi\circ\phi^{{\dagger}}$ is an auto-encoder, which can encode the input data $X$ into a latent space $U$ and reconstruct $X$ according to $U$. While, according to Equation 51, the latent variables are independent each other. That is, we represent the input data by a set of independent or disentanglement latent variables, and we can reconstruct the input data according to these independent variables. Therefore, this is called disentanglement representation learning. In practice, independent component decomposition method can be use to construct the decoder Hyvarinen and Morioka (2016); Khemakhem et al. (2020). To reduce the computational complexity, we can also use invertible neural network to model $\phi$ Dinh et al. (2014).
With the disentangled representation variables $U_{i}$, we can effectively perform downstream tasks. One prominent application of disentangled representation learning is causal discovery. By leveraging the disentangled representations, we can uncover causal relationships between variables and gain insights into the underlying causal structure of a system Monti et al. (2020). Causal discovery allows us to identify cause-and-effect relationships and understand how different factors interact and influence each other. This knowledge can be invaluable in various domains, such as understanding complex systems, making informed decisions, and designing interventions or policies.
To conclude, a typical causal machine learning or disentanglement representation learning framework is always composed of a module for causal or disentanglement representations and a module which contains sub-units for different downstream tasks. The representation module, which contains an encoder and a decoder, tries to extract abstract representations of the data, and the representations are not for prediction but for causal inference. The downstream tasks may be variant. For example, the transferable mechanism learning tries to look for independent causal mechanisms to complete prediction tasks in very different environments Peters et al. (2016). And existing studies have shown that learning causal models may help in transferring modules across different domains Cui and Athey (2022); Kuang et al. (2018). Causal representation learning has been applied to various problems.
4.2.2 Comparisons between Causal Emergence Identification and Causal Representation Learning
There are similarities and differences between causal emergence and causal representation learning (CRL). We will introduce them in detail separately.
Similarities
It is interesting to compare the tasks of causal emergence identification under the theoretic framework of Erik Hoel and causal representation learning. Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro variables are the causal variables, the macro dynamics are the causal mechanisms, the coarse-graining strategy is an encoding process from the original data to the representations. Under this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism. To identify causal emergence, we need to learn the appropriate causal variables to represent data and to discover causal mechanism on the macro-level (on representational latent space) such that the EI of the learned causal mechanism in latent space is larger than the original data. We use Table 1 to clarify the comparisons item by item.
Table 1: Comparison between causal emergence identification and causal representation learning.
Causal Emergence Identification Causal Representation Learning
Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real
Latent Variables Macro-states Causal representations
Causal mechanism Macro-dynamics Causal mechanisms
Mapping between data and latent variables Coarse-graining Representation
Optimization for causality EI maximization Prediction Loss, Disentanglement
Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has stronger causal effect Finding an optimal representation of the raw data such that the independent causal mechanism can be realized by the representations.
With these similarities, the technologies and concepts from both fields can learn from each other. For example, the techniques of causal representation learning can be applied to discover causal emergence.
On the other hand, the learned abstract causal representations can be viewed as macro-states, which enhances the explainability of causal representation learning. This perspective allows us to consider that CRL is actually uncovering the hidden causal emergent features within the data.
Moreover, these similarities between emergent phenomena and CRL contribute to a deeper understanding of emergence itself. By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract a more profound causal structures. These profound causal structures may serve as the origins of the mysteries surrounding emergence. By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we gain insights into the mechanisms behind emergent phenomena. These causal structures provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components. Exploring these intricate causal structures brings us closer to unraveling the enigma of emergence and sheds light on the fundamental principles governing complex systems. Consequently, the identification of emergent phenomena requires an observer, whether it be a human or a machine, to actively search for them.
By employing a machine observer, it is possible to potentially resolve the longstanding debates regarding whether emergent causality is epistemic or ontological. This is because the machine observer can provide a relatively objective outcome, although it is important to note that the result is still influenced by the observer. Furthermore, an intriguing area for future exploration involves investigating how the structure and capabilities of the observer can impact the emergence results that are identified by modifying the parameters of machine learning. This highlights the potential influence of the observer’s characteristics on the outcomes obtained through the rewiring of machine learning parameters.
Differences
However, there are several differences between the CRL and causal emergence in theory. The biggest gap is that CRL assumes that there is a real causal mechanism behind it, and the data is generated by this causal mechanism Schölkopf et al. (2021). While for causal emergence identification, the emergent variables and mechanisms on macro-level are just handy ways to observe and understand data, "real causality" may not exist. The difference can be illustrated by Figure 17
Figure 17: The Comparison between Causal Emergence Identification and Causal Representation Learning. All the potential co-founders are ignored. The black arrows represent causal relations, while the blue arrows represent interactions between macro-variables. The dashed directed lines represent the information flows for coarse-graining, representation or data generation.
While, if epistemological perspective is adopted, this difference disappears because all they do is to extract meaningful information from observational data to obtain representations with stronger causal effects.
Another major difference is that the causal mechanism for macro-states in causal emergence is a dynamical system. Therefore, if the dynamics in macro-state space is on a network, circular structures may exist because the state variables will be iterated in the dynamical system. As a result, feed-backs and circular indirect interactions are allowed in such model. However, the causal mechanism is always represented by a structural causal model in CRL. While, these models are always directed acyclic graphs, therefore, loops or circular structures are not allowed. The problem is that the meaning of the links in a network dynamics model is different from the links in structural causal model. Hence, the conflict can be resolved if we comprehend that the links in a structural causal model depict an average (net) causal effect between variables when they are in a steady state. On the other hand, the links in network dynamical models within causal emergence represent instantaneous interactions.
Finally, unknown co-founders are always ignored in causal emergence, while, they play very important roles in causal structural models.
4.2.3 Out-of-distribution Generalization
Causal representation learning is established in a self-supervised manner, in the sense that it reveals the intrinsic causal structure of data. Thus, the learned causal representations could not be exploited directly for supervised prediction task, where the causation between data and a specified target must be characterized. The measurement of causal emergence by vanilla Effective Information Hoel et al. (2013) is also defined for stationary dynamics systems in a self-supervised fashion, since the input(cause) space and the output(effect) space share the support. However, Hoel generalizes Effective Information to general input-output systems Chvykov and Hoel (2020)(see section 3.6.1 where the cause and the effect space take arbitrary support, which could either be discrete or continuous.
Empowered by the general notation of EI, causal emergence could be applied in supervised learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or a representation learned upon which) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause(feature) to the effect(target). It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffice for common prediction tasks with an i.i.d. assumption, implying that the training and test data are independently and identically distributed.
However, if samples are drawn out of the training distribution, it is essential to learn a representation space that generalizes from training to test environment, known as out-of-distribution (OOD) generalization. Due to the common belief that causal relations generalize better than statistical correlation Arjovsky et al. (2019), causal emergence theory could serve as a criteria for the causation embedded in the representation space. The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regarding to out-of-distribution generalization.
Introduction of OOD
Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic. Out-of-distribution generalization towards arbitrary distributional shifts is generally impossible Ye et al. (2021). Therefore, it is commonly assumed there exists an underlying causal mechanism that governs the data generation process, which remains invariant to distributional shifts. Formally, by Pearl’s Structural Equation Model Pearl (2009) which causally formulates data generation, the target $Y$ is supposed to be generated from part of the feature $X$, i.e. $Y=f(X_{c},\epsilon)$ where $X_{c}$ is selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$. Obviously $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments. On the other hand, the raw feature $X$ might be a mixture of causal variables and non-causal variables. The latter could be produced by confounding, selection bias or anti-causal mechanisms, all of which introduce spurious correlation between the target and non-causal variables. The strength and direction of spurious correlation are variant across environments, resulting in distributional shifts that hampers out-of-distribution performance of machine learning techniques. As a remedy, a prominent branch of OOD algorithms Arjovsky et al. (2019); Huh and Baidya (2022); Zhao et al. (2019); Ganin et al. (2016) learn an invariant representation space distilling the causal variables from the raw feature space. Typically a metric of the representation is devised to verify the stability of correlation between the learned representation space and the target, and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features. The paradigm for invariant representation learning could be formulated as Wang et al. (2022):
$\min_{f,\phi}\mathcal{R}(f(\phi(X)),Y)+\lambda\ell_{reg}(\phi),$ (54)
where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function and $\ell_{reg}$ is the regularization.
Representation learning based OOD algorithms are diversified in their constraint of the representation space characterized by the regularizers. Each metric of the representation fits a specific data generation pattern, and does not guarantee to filter out the causal variables when the setting is violated. However, causal emergence potentially offers a united metric of representation for OOD generalization rooted from causality theory.
Before diving into the application of causal emergence, we would first discuss representative regularizers of representation learning and their limitations. Domain invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments. The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al. (2018) or an adversarial classifier for the environment index Ganin et al. (2016). However, the alignment constraint provides generalization guarantee only for covariate shift Zhao et al. (2019) where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$. On the other hand, invariant risk minimization Arjovsky et al. (2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$. Invariant risk minimization is proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target, but is reported to fail under covariate shift Ahuja et al. (2021). Further, conditional invariant representation Li et al. (2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect. The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
Actually, it is theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models and it does not generalize across all data generation processes Kaur et al. (2022). As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, which hinders the general applicability of invariant representation learning. Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedure, increasing the data collection and labelling expenses. To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
Causal emergence and OOD
The match between OOD generalization and causal emergence is attributed to the do intervention of Effective Information. By definition Pearl (2009), intervening the feature distribution with $do(X)$ modifies the structural equations of all the variables in feature space to a constant value while the equation for the target remains intact as $Y=f(X_{c})$. On the one hand, the assumption of causal mechanism’s invariance to distributional shift is respected, since the causal correlation between causal variables and the target is reserved. On the other hand, the spurious correlation between non-causal variables and the target are removed by the intervention. For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation by $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$. As a result, $X_{n}$ is no longer caused by $Y$ and the spurious correlation vanishes. For another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$. Then the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist. In summary, the intervention introduced by EI captures the causal correlation between features and the target whereas suppresses the spurious correlation. It renders EI an ideal metric to assess the causality contained by the representation space and thus the generalization of the induced algorithm.
Recall that an increment in the Effective Information implies the occurrence of causal emergence. Therefore, we bring forward a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning. Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e. $I(g(X);Y)\leq I(X;Y)$, by data processing theorem. Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space. In contrast, EI(or normalized EI) would possibly peak at a medium stage of abstraction of raw feature, coinciding with the philosophy of OOD generalization that less could be more. Ideally, at the peak of EI(or normalized EI) where causal emergence occurs, all non-causal features are excluded and the causal features are revealed, resulting in the most informative representation while performing invariant to distributional shifts.
Actually, removing spurious correlation and recovering causal mechanisms by manipulating feature distributions is shared by the reweighting technique widely adopted in OOD literature. For instance, Stable Learning Shen et al. (2020a, b); Zhang et al. (2021) is oriented at the setting where collinearity exists in input variables, causing spurious correlation between non-causal variables and target. It reweights training samples to decorrelate the features and thus reducing collinearity and eliminating spurious correlation. Similar to EI, Stable Learning is free of environment index labeling. Further, sample reweighting and feature decorrelation shares the philosophy of distribution intervention. In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
4.3 Other Related Topics in AI
There are other related topics in artificial intelligence, e.g., abstraction of causal models and world model based reinforcement learning. We will introduce them in detail.
4.3.1 Abstraction of Causal Models
In CRL, we have the ability to extract causal relationships and construct causal models from raw data. However, in cases where the causal model becomes excessively large and complex, it becomes necessary to employ an abstraction process Rubenstein et al. (2017). Abstraction in causal modeling refers to the reduction or simplification of the causal model while preserving its essential causal relationships and important features.
Coarse-graining is the multi-scale description of the state information on different information reduction scales. While causal models also can be multi-scale Chalupka et al. (2016). The causal model in the higher level is an abstraction of the model in the lower level.
If the original causal model is given, then the key problem is how can we obtain an abstraction of the original model. Obviously, the abstraction of the causal model cannot be arbitrary, it must recover the main features of the original causal model. Several works proposed different standards for abstract models Rubenstein et al. (2017); Beckers and Halpern (2019); Chalupka et al. (2016), among which approximate abstraction is a good and reasonable proposal. The basic idea behind the method of approximation abstraction is that the differences of causal mechanism both in the abstraction and the original causal model should be minimized Rubenstein et al. (2017); Beckers and Halpern (2019); Beckers et al. (2020); Beckers (2021).
Figure 18: An approximate abstraction of the El Niño phenomenon.
For example, we can build a causal model to describe how the Equatorial Pacific Zonal wind ${W}$ can have a causal effect on the temperature of the Equatorial Pacific ${T}$ Chalupka et al. (2016). We write the causal variable $W$ as $X_{C}$, the effect variable $T$ as $X_{E}$, and this is the original causal model, which can be denoted as $\mathcal{M}_{m}=((\{{U}\},\{{X_{C}},{X_{E}}\},\mathcal{R}_{m}),\{g_{m},f_{m}\},\mathcal{I}_{m})$, where ${U}$ is the collection of all the unknown variables, ${X_{C}}$ and ${X_{E}}$ are causal variables, $g_{m}$ and $f_{m}$ are causal mechanisms to describe how ${U}$ and ${X_{C}}$ can affect other variables, $\mathcal{I}_{m}$ is the set of interventions. However, we care about higher level causality when we consider the El Nino or La Nina phenomena. Thus, we can build an abstract causal model $\mathcal{M}_{M}=((\{{U}\},\{Y_{C},Y_{E}\},\mathcal{R}_{M}),\{g_{M},f_{M}\},\mathcal{I}_{M})$, where $Y_{C},Y_{E}$ are new causal variables which are built by the coarse-graining mapping $\phi$, that is $Y_{C}=\phi({X_{C}})$, $Y_{E}=\phi({X_{E}})$. And $g_{M}$ and $f_{M}$ are causal mechanisms of $Y_{C},Y_{E}$ on the abstraction model. However, if $\phi$ is given, $g_{M}$ and $f_{M}$ are limited otherwise inconsistencies will exist. In  Beckers et al. (2020), the authors propose a limitation on $g_{M}$ and $f_{M}$. We denote $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ as the distance between different level models $\mathcal{M}_{m}$ and $\mathcal{M}_{M}$. They claim that $\mathcal{M}_{M}$ is an $\epsilon$-approximate abstraction of $\mathcal{M}_{m}$, if and only if $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ must be less than $\epsilon$ Beckers et al. (2020). That is,
$d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})=\min_{\phi_{\mathcal{U}}}\max_{do(X=x)\in\mathcal{I}_{m},{u}_{m}\in\mathcal{R}_{m}(\mathcal{U}_{m})}$
$\big{\{}d_{v_{M}}(\phi(\mathcal{M}_{m}({u}_{m},do(X=x))),\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x))))\big{\}}.$ (55)
In this equation, $\phi_{\mathcal{U}}$ is a coarse-grained mapping of unknown exogenous variables, $do(X=x)$ is an intervention on certain endogenous variables, $d_{v_{M}}$ is the distance between variables defined at a higher level, and $\omega_{\phi}$ is the mapping of interventions between different causal models at different levels. The distance between models is transformed into the distance between variables. Given a defined method of measuring the distance between variables (such as Euclidean distance), we can measure the distance between the output of a lower level model after coarse-grained projection$\big{(}\phi(\mathcal{M}_{m}({u}_{m},do(X=x)))\big{)}$ and the variables output by a higher level model$\big{(}\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x)))\big{)}$. We focus on identifying external variables and intervention methods that maximize the distance, and we do not concern ourselves with the coarse-grained mapping between external variables. Therefore, we select the optimal $\phi_{\mathcal{U}}$ and require that it be surjective Beckers et al. (2020). This yields a distance value that characterizes the inconsistency between different levels of causal models.
There are similarities between causal emergence identification problem and abstraction of causal models. If the original causal mechanism can be understood as a micro-dynamics, then the abstracted mechanism can be regarded as a macro-dynamics. Therefore, finding the optimal $\phi$ and causal mechanisms $g_{M}$ and $f_{M}$ is equivalent to finding an optimal coarse-graining strategy and macro-dynamics. In the framework of NIS, we set limitations on coarse-graining strategy and macro-dynamics by requiring that the macro-dynamics can recover the micro-dynamics with $\epsilon$ error to exclude trivial cases. This requirement is also similar to the approximation requirement for the distance.
In artificial intelligence, model abstraction is a widely used technique to obtain a compressed model. For example, the abstraction of a neural network model can be stored in cell-phone to conduct fast computation Wang et al. (2021); Pulina and Tacchella (2010). Another article discusses the importance of structural analysis methods for neural network analysis and proposes a new method based on a formal theory of causal abstraction. This method aligns neural representations with variables in interpretable causal models and uses interchange interventions to experimentally verify their causal properties. The method is utilized in a case study to examine neural models that have been trained on a complex Natural Language Inference (NLI) dataset Geiger et al. (2021). This dataset comprises numerous artificially processed, linguistically intricate English phrases and sentences. The analysis reveals that neural representations effectively capture the compositional structure of the provided examples Geiger et al. (2021). Model reduction is also important topics in control theory and dynamical systems Antoulas (2005).
The primary distinction between causal emergence identification and causal model abstraction lies in their approach to abstraction or dimension reduction. Causal emergence identification focuses on abstracting or reducing dimensions from the original data, whereas causal model abstraction focuses on abstracting from the micro-level causal models. In another word, coarse-graining operates on states or data, while abstractions operate on models. In physics literatures Christensen and Moloney (2005), abstractions of models are also called renormalization.
Another difference is unknown confounder $\mathcal{U}$ being considered in causal model abstraction is ignored in all the discussion of causal emergence. The emergence of causality requires the construction of models from a multi-scale perspective, and how to discuss mixed variables at different scales has always been a challenge. The theory of approximate abstraction takes into account the coarse-grained mapping of latent confounding variables $\phi_{\mathcal{U}}$, making it compatible with the coarse-grained function of endogenous variables. For the study of causal emergence, this theory lays a theoretical foundation for considering the confounding effects brought by environmental factors.
The approximate abstract framework currently has many limitations. Firstly, there is a lack of effective algorithms that can calculate the degree of approximation in the approximate abstract. Given the involvement of traversing intervention sets, it is difficult for machines to perform efficient numerical calculations that are not maintainable, and cannot be used as constraint conditions for automatic modeling. Additionally, the intervention space is very large, and some interventions are meaningless or impossible, so it is necessary to empirically provide an intervention distribution in advance Beckers et al. (2020), making it difficult for this framework to be applied to real-world problems.
4.3.2 World Models in Reinforcement Learning
In both the generalized causal emergence identification problem and causal representation learning, a common computational framework is shared, consisting of an encoder-decoder architecture and causal mechanisms or dynamics. This framework can also be integrated into a reinforcement learning agent, enabling the agent to learn a world model and make informed decisions and plans. Such an agent, known as a world model-based reinforcement learning agent, has demonstrated impressive performance across various environments and tasks Ha and Schmidhuber (2018).
Reinforcement learning is a significant sub-field of machine learning that possesses several distinctive features when compared to conventional supervised learning techniques Sutton and Barto (2018). Firstly, reinforcement learning agents have the capability to interact with the environment and actively acquire new information through interventions. Consequently, causal information can be obtained through these interactions Zeng et al. (2023). Secondly, in contrast to supervised learning where the processes of learning and action are clearly separated, reinforcement learning agents can seamlessly blend learning and action processes. Lastly, reinforcement learning agents demonstrate strong performance in environments that have delayed feedback, such as game playing scenarios Mnih et al. (2013); Silver et al. (2016).
The reinforcement learning algorithms can be classified into model-based and model-free types, and the world model based reinforcement learning belongs to the former Ha and Schmidhuber (2018); Hafner et al. (2023); LeCun (2022); Hafner et al. (2019a, b, 2020, 2023). As its name suggests, there is a model of the world which can simulate the dynamics of the environment the agent confronts. The dynamics of the world model can be learned via the interactions between the agent and the environment, and it can also help the agent to make plans and decisions about the uncertain environment Sekar et al. (2020). However, in order to represent a complex environment, the world model needs to provide a coarse-grained description of the environment. Consequently, a standard architecture of a world model typically includes an encoder and a decoder, as illustrated in Figure 19.
Figure 19: The architecture of a world model embedded in a model-based reinforcement learning agent, where $X_{t}$ is the environment at time $t$, and $Y_{t}$ is its representation, $H_{t}$ is the internal variable of the agent. Different from NIS and causal representation learning agent, another variable, the action in the previous step $A_{t-1}$, should be input to the world model to implement the one step dynamics. And the world model can be trained in a supervised learning manner by comparing the predicted representation of the environment and the prediction.
The world model can also be treated as a macro-dynamics or causal mechanism Zhang et al. (2020); Allshire et al. (2021); Jiang et al. (2022); Zhang et al. (2020); Wang et al. (2022). All the states of the environment can be regarded as macro-states or causal representations. These states are constructed to ignore irrelevant information and capture the most important causal features of the environment for the agent to make decisions. In the process of planning, the agent can use the world model to simulate the running of the real world dynamics. Numeric results show that the reinforcement learning algorithms with world models can perform better than the classic reinforcement learning algorithms.
World model-based reinforcement learning techniques can also be employed for identifying causal emergence. When attempting to identify the causal emergence phenomenon within a highly complex system, extracting macro states and dynamics solely from observational data becomes challenging. Explorations or experiments are required to facilitate this process Sekar et al. (2020). In such cases, a reinforcement learning agent equipped with a trainable world model, specifically a macro-dynamics model, becomes essential. By treating the complex system as a unified entity, the agent with the world model can interact with it and derive emergent causal laws from these interactions.
On the other hand, the exploration of causal emergence can offer valuable insights into reinforcement learning algorithms. Firstly, the measure of causal effect, EI, can be applied to world models. Techniques of maximizing EI can be employed to enhance the causal effect of world models within reinforcement learning agents Yang et al. (2023). This, in turn, can greatly benefit planning and problem-solving tasks, improving their efficacy and performance. It is like the bi-simulation Kemertas and Aumentado-Armstrong (2021); Zhang et al. (2020) which can train the agent with an internal metric.
Second, if the reinforcement learning agent with a world model can learn the emergent causal mechanisms automatically from a complex system that the agent is interacting with, then it will have more complex behaviors and be more intelligent. The agent may use the learnt emergent causal mechanisms to make plans. From learning to planning, the whole process may take place continuously. Therefore, causal emergence can be observed from outside because new behaviors governed by different causal mechanisms emerge as time goes by. This agent seems to have its own purpose or objectives.
4.4 Summary
In summary, there are deep connections between causal emergence and some hot topics in artificial intelligence such as CRL, abstractions, and world models. On one hand, the machine learning framework for causal emergence identification has similar structures to the frameworks of CRL. And even the concepts can have one to one correspondence as shown in Table 1. Therefore, the ideas and techniques of both sides can learn from each other. On the other hand, the studies of causal emergence can provide new perspectives and insights for causal representations and world model based algorithms. For example, we can explain that what the representations and causal mechanisms learned are emergent causal laws. Maximization of EI may help to improve the efficiency of learning the causalities for both CRL and reinforcement learning agents. Finally, we found a reinforcement learning agent with the ability to learn causal emergence may exhibit the ability of actions for purpose and objectives.
5 Potential Applications in Complex Systems
In this section, we will explore how the problems discussed in the previous sections relate to our understanding of complex systems. On one hand, the methods of causal emergence identification and the associated AI algorithms can be employed in complex systems to discover emergent patterns and causal relationships. On the other hand, the concepts of causal emergence can aid in our comprehension of complex systems, particularly in addressing significant issues such as consciousness and free will. These ideas have the potential to provide valuable insights and approaches to tackle these complex and profound problems.
5.1 Potential Applications
Complex systems typically consist of a large number of interacting units, and these units and interactions form various hierarchical and self-similar structures in a complex manner Mandelbrot and Mandelbrot (1982). In particular, when we examine complex systems from different scales, various structures and patterns may emerge at different scales, and these patterns may further interact with each other. Thus, a complex network of causality may form, and the nodes in this network may not only include the individual units of the complex system but also contains the emergent patterns in distinct scales.
As a result, it is crucial to determine how we can identify and extract emergent patterns and causal relationships from vast amounts of data generated by complex systems.
5.1.1 Emergence detection
In complex dynamical systems, various higher-level patterns, such as waves, periodic oscillations, and solitons, are ubiquitous. For instance, in the climate system, typhoons and tornadoes are emergent vortex structures. In urban areas, traffic jams also emerge as a result of interactions among a large number of cars. Social riots are another example of emergent events in higher levels of human society. Identifying these emergent patterns at an early stage is crucial and significant Haugen et al. (2023); Mnif and Müller-Schloer (2011). Therefore, there is an urgent need for a method that can automatically detect these emergent patterns and even provide early warning signals.
In O’toole et al. (2017), the authors regard emergent phenomenon as a specific event happened in a complex system, and detecting emergence is equivalent to detecting the formation and evaporation of this event. Therefore, emergence detection problem can be converted as a change-point detection problem of time series data.
Two methods are discriminated to detect emergence: one is centralized method, the other is decentralized. For example, in Mnif and Müller-Schloer (2011), the authors propose that the negative Shannon entropy can be used to measure the orderness of the entire system, and it can also indicate if emergence occurs. NIS and NIS+ are also centralized methods.
Another one is decentralized emergence detection method, in which the detectors are dispersed on the complex system, and implement event detection spontaneously and separately O’Toole et al. (2014).
As the development of artificial intelligence, some deep learning methods can be used to detect emergence. For example, dynamical image segmentation can be used to find interesting patterns in the video automatically Liu et al. (2021). A complex system can always be represented as a dynamical graph, e.g., a flock of birds, and when emergence occurs, the connectivity of this interaction network always increase. Therefore, we can detect the change of the network to find emergence Pazho et al. (2022). Other deep learning methods could be helpful.
5.1.2 Discovering emergent causality
Complex systems can be understood as a large network with events connected by causal links. A small disturb of one event may give rise a large roam of changes as butterfly effect indicates. While, sometimes, one single event happens with destiny which means it may be the effect of many different causes. How to discover the complex causal relationships from data in an automatic way is a challenging problem.
The research area of causal discovery tries to tackle this problem in many different ways Glymour et al. (2019); Zanga et al. (2022). For example, some studies try to use well-defined statistical indicators to characterize the causality between two variables, e.g., the measures of mutual information Donges et al. (2009), transfer entropy Stetter et al. (2012), etc. Nevertheless, these methods can not distinguish causal relationship from correlations. Another widely used method is independent causal discovery method Glymour et al. (2019), which tries to find the conditional independence among three variables in data, and derive Markov equivalent sub-graphs accordingly. Finally, binary causal interactions can be identified from these sub-graphs Kalisch and Bühlman (2007). Score based method tries to maximize or minimize a score function to find the optimal causal structures Zheng et al. (2018); Zhu et al. (2019). These methods always require a large amount of data and computation.
However, more challenges will be confronted when these causal discovery methods are applied in complex systems because circular causal structures and cross-level causality exist.
Circular causal structures widely exist in complex dynamical systems due to the feedback effect. That is, one variable may take effect on itself via a direct self-loop feedback or a circle with long chain. This circular causal structure may challenge the existed study methods of causality because most of these methods study directed acyclic causal graphs. Recent progress has developed some methods to discover these circular structures e.g.  Richardson (2013); Lacerda et al. (2012); Forré and Mooij (2018); Rantanen et al. (2020); Zhang et al. (2019); Pamfil et al. (2020); Vowels et al. (2022).
Another aspect is higher level or cross-level causation if the ingredients of scale and coarse-graining are considered. For example, downward causation describes the causal effect between macro-level and micro-level variables. Thus, the causal connections may be cross-leveled. If causal emergence occurs in a complex system, some strong causality may be found between macro-variables. All of these factors in complex systems present new challenges. The machine learning methods discussed in section 4.1.2 can address the question of whether there exists higher-level causality at the macro-level. However, it’s important to note that this is a global property. It must be further developed and extended to find local causal relationship between macro-variables or macro-variable and micro-variable Varley (2022). While, existed methods of causal discovery must be extended to take the method of grouping a set of variables or coarse-graining a system into consideration. For example, the article Cheng et al. (2023) proposes a coarse-to-fine causal discovery algorithm based on Granger causality and graph neural network. The grouping process of variables is from coarse-to-find which can improve the efficiency of the algorithm. Other multi-level causal discovery methods are also proposed and discussed in Wang et al. (2023); Zhang et al. (2023); Fan et al. (2023).
One of the potential applications of these complicated causal discovery method is to enhance the interpretabiity of large and complex neural networks. As we know, interpretability is one of the notorious problem for large neural networks, although they are very useful to solve many real problems. Thus, we can apply the methods of causal discovery, especially, the methods to discover emergent causality on these neural network models to find hidden and emergent causal relationships between neurons or group of neurons Chattopadhyay et al. (2019); Ren et al. (2023). If the degrees of causal effect can be computed and compared among these discovered causal relationships, the most important causation can be pointed out to solve a concrete problem. This may enhance the interpretability of the neural network. In this way, the misery of emergent ability of large language models may be dismissed by uncovering the hidden causation within large models Michaud et al. (2023); Wei et al. (2022).
5.1.3 Beyond causality
In this review, we mainly focus on the emergence of causality, however, we do not need to be bounded on this kind of special emergence, more macroscopic or global level properties could also be emergent although causal emergence may be the most important one. For example, individuality is one of the emergent property with no any connection with causality which has been discussed in previous section. In  Chang et al. (2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world. The authors then argued that this agent with information closure be regarded as a model of consciousness.
Symmetry may be another interesting property within some complex systems, and it may be emergent. For example, any single trajectory of a large number of random walkers in a two dimensional Euclidean space is random, nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic. This kind of symmetry can only be found on the macro-level, thus, is emergent.
It should be noted that EI can also be viewed as a unique indicator of symmetry, specifically, it is time reversal symmetry. The reason behind this is that when the macro-dynamics approaches a time reversible map (a bijection between states across a time step), EI is increased. And the maximum EI can be arrived when the Markov dynamics is reversible. This interesting fact should be further studied in the future.
It is interesting and useful to deploy a machine learning method to automatically discover this kind of emergent symmetry. The framework of NIS may be extended to account for the problem of emergent symmetry identification if the optimized objective, EI, is replaced with some measure of isotope or more general symmetry Liu and Tegmark (2022). Similarly, we can also find other emergent symmetry with this framework once the symmetry can be measured and optimized.
5.1.4 Designing Emergence
We even can design a set of interaction rules for a complex system such that the desired property can emerge from the interactions once the desired property can be defined in a clear and quantitative way. The basic idea to realize emergence design is by optimization Adamatzky et al. ; Kreyssig and Dittrich (2011). Actually, the early studies of evolutionary computation or genetic algorithm have solved this problem for some simple and small cases Adamatzky et al. . For example, M. Mitchell et al. used genetic algorithm to search the interaction rules of one dimensional cellular automata such that the automata can solve some simple mathematical problems Mitchell et al. (1993). Recently, as the development of machine learning and automatic differential, searching such interaction rules may become easier and more efficient if the rules can be converted into differentiable operations Mordvintsev et al. (2022); Gala et al. (2023); Zhang and Walker (2022).
However, if the space of possible interaction rules is very large, all the methods based on search or optimization may be invalid. Thus, some ideas developed in the field of deep learning can be borrowed here to design emergence. For example, the framework of generative adversarial network (GAN) can be used to design complex systems to realize some emergent properties Goodfellow et al. (2014). On one side, we should design a generative model for a complex system or the interaction rules for the system such that a large number of candidate rules or systems can be generated. On the other side, a discriminator or a critic can be designed to evaluate the candidates if the desired emergent properties have been realized. Similar ideas can be used on other possible generative models, e.g., diffusion model, normalizing flow Cao et al. (2022); Yang et al. (2022); Papamakarios et al. (2021); Kobyzev et al. (2020), etc.
However, the difficulty of these methods lies in the demand for a large amount of training data and computations. The lack of data may hamper the applications. In general, designing emergence is a useful problem but it is also difficult.
5.2 Understanding Complex Systems from Causal Emergence
Deep understanding of causal emergence and emergent causality will help us to understand the hidden mechanism of complex systems including living systems, social systems, climate systems, ecosystems, etc. Particularly, some big problems like free will, consciousness Tononi and Koch (2015), and life are all related to emergent causation Walker (2014). For example, free will can be understood as an emergent downward causation Yurchenko (2023). In Chinese medicine, Yin and Yang can be understood as macro-states and the properties of their interactions are emergent causality. Social phenomena can be understood from the viewpoint of causal emergence Elder-Vass (2010). For instance, the increase of the rate of suicide may be the emergent causal effect of economic crises. Interestingly, both EI and $\phi$ID frameworks on causal emergence have relation with one of the consciousness theory: integrated information theory Tononi et al. (2016). As a result, further studies of causal emergence and the connections with other big problems are required.
For a particular system, there are three problems that should be addressed in future studies: 1. When does causal emergence occurs? 2. How emergent causality take functional effects on the system? 3. How emergent causality may change when the system is changed to adapt the environment? We will discuss these problems one by one in brief.
We still do not know when will causal emergence occurs and how does the measure of causal emergence change with some key parameters of a system. In Yang et al. (2023), the authors have shown how the measure of causal emergence change with different noises in an extended Boid model. It increases as the observational noise increases because coarse-graining operation can eliminate the observational noise. However, the same coarse-graining operation can hardly reduce the intrinsic noise (the noise within the dynamical equations), which leads to the degree of causal emergence decreases. It is reasonable to expect that there is a phase transition of the emergent causality within a complex system when some key parameter changes because the causality or the strength of causal effect is also a global property and it may dependent on some order parameters.
If causal emergence occurs in the system, how does the emergent causation take effects on the part and the whole of the system? For example, the emergent "I" can be understood as an emergent macro-variable Yurchenko (2023). How does "I" can influence the part, say my foot, to move? This problem is non-trivial because it relates to the problem of mind-body interaction. One of the difficulty is that how can we build a model for such systems? Should we directly model the emergent properties and causation or model the underlying interaction rules and ignore the causation in higher level? If we directly model the emergent "I", the model may become very complex because the properties of "I" are difficult. However, if we model the system from the micro-level interactions, we can still hardly understand the emergent causality. Shall we switch the viewpoints between micro- and macro-? If yes, in what manner to switch?
Finally, what is the relationship between adaptation and emergence Holland (1995, 2000, 1992)? Sometimes, when we refer to a property as emergent, we actually mean that this property can be developed through adaptation. Therefore, adaptation or evolution serves as a causal force for certain emergent properties. This concept also applies to emergent causality. For instance, downward causation, which is commonly found in complex systems, emerges as a result of adaptation and evolution. The next crucial question is: how can we evolve an emergent property or causation? This problem resembles the issue of designing emergence discussed in Section  5.1.4. However, our aim here is to seek an explanation rather than a design. We want to understand the specific environment and manner in which the observed emergent causation can be evolved through adaptation.
All these problems desire further studies in future. However, these problems are just name a few, more interesting problems need to be addressed.
6 Conclusion
Quantifying the notion of emergence has been a challenge for a long time. Causal emergence serves as an important step in the development of a quantitative theory of emergence, making it a crucial concept in the field of complex system studies. In this paper, we provide a systematic review of papers on causal emergence and other related quantitative studies pertaining to emergence and causality. Our analysis highlights two crucial problem areas: emergence quantification and causal emergence identification in data. We discovered intriguing connections between the latter problem and artificial intelligence, particularly in sub-fields such as causal representation learning and world-model based reinforcement learning. As artificial intelligence continues to advance, we anticipate that more puzzles related to causality and cross-level interactions within complex systems will be solved.
7 Appendix A - Calculation of EI with TPM
Calculating effective information ($EI$) requires knowledge of the joint probability distribution of the system. The state transition probability matrix (TPM) defines the probability of transitioning from one state of the system to another. Specifically, the $j_{th}$ element of the $i_{th}$ row in the TPM, denoted by $TPM(i,j)$, represents the probability of the system being in state $s_{j}$ at time $t+1$, given that it is in state $s_{i}$ at time $t$ (i.e., $P(X_{t+1}=s_{j}|X_{t}=s_{i}))$.
$TPM(i,j)=P(X_{t+1}=s_{j}|X_{t}=s_{i})\\ =\frac{P(X_{t+1}=s_{j},X_{t}=s_{i})}{P(X_{t}=s_{i})}$ (56)
Assuming we can "force" the system to conform to a maximum entropy distribution (i.e., uniform distribution) with $n$ states, the probability of being in any given state $i$ at time $t$, denoted by $P_{u}(X_{t}=s_{i})$, is $1\over{n}$.
As a result, the TPM can be expressed as the product of the joint probability of the system’s maximum entropy distribution, denoted by $P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})$, and the total number of states, which is $n$, as demonstrated below.
$TPM(i,j)=\frac{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}{P_{u}(X_{t}=s_{i})}\\ =\frac{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}{1/n}\\ =n\times{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}$ (57)
Hence, the TPM of the system can be used to express both the joint probability and marginal probability at time $t+1$ of the system’s maximum entropy distribution.
[…]
Using this relationship, $EI$ can be calculated directly with the TPM.
[…]
Reference